### v1

# Relativistic framework for high-precision GNSS-based navigation in cislunar space

Abstract
[…]
I Introduction
High‐precision spacecraft navigation requires a fully relativistic treatment of both time and space observables [1, 2, 3]. Central to this treatment is the choice of an appropriate coordinate reference system. The Barycentric Celestial Reference System (BCRS), with origin at the solar‐system barycenter, underlies deep‐space applications such as lunar laser ranging (LLR), very long baseline interferometry (VLBI), and the construction of planetary and lunar ephemerides [4]. By contrast, the Geocentric Celestial Reference System (GCRS), centered on Earth’s mass center, is the natural frame for precise orbit determination of GNSS and other Earth‐orbiting satellites [2]. Transforming observables between these frames requires inclusion of all $O(c^{-2})$ corrections—gravitational time dilation, velocity‐dependent terms, multipolar potential perturbations, and frame‐dragging effects—to avoid systematic errors in both timing and positioning.
For Earth-orbiting spacecraft, GCRS-based models incorporate relativistic corrections to the Newtonian equations of motion, including adjustments to central gravitational acceleration and terms arising from Lense-Thirring precession (caused by Earth’s angular momentum), geodesic precession (due to spacetime curvature), and Coriolis effects [5, 6]. These relativistic corrections, while relatively small compared to non-relativistic forces such as atmospheric drag and solar radiation pressure, are critical to achieving the cm-level positional accuracy required for modern navigation systems. Over time, accumulated relativistic errors could cause significant positioning drift in long-duration satellite missions, making it essential to integrate these terms into orbit determination models to prevent systematic errors.
For interplanetary and lunar trajectories, spacecraft dynamics are governed by the Einstein–Infeld–Hoffmann equations in the BCRS [7, 1]. The differential rate between Terrestrial Time (TT) and Barycentric Coordinate Time (TCB), $1.5505\times 10^{-8}\,\mathrm{s/s}$, yields $\sim 1.34\,\mathrm{ms/day}$ offsets near Earth, significantly affecting one‐ and two‐way light‐time calculations [8, 9]. In strong‐field regions—e.g. at ten solar radii where the Parker Solar Probe operates—the relativistic correction to Newtonian acceleration ($\sim 1.22\times 10^{-6}\,\mathrm{m/s^{2}}$) becomes non‐negligible relative to the $2.74\,\mathrm{m/s^{2}}$ background [10, 11]. Even at 1 AU, frame‐dragging and potential‐induced time dilation ($\sim 9.87\times 10^{-9}\,\mathrm{s/s}$) must be applied when transforming between geocentric and barycentric frames to maintain picosecond‐level timing consistency.
As GNSS technology advances with the goal of achieving cm-level positional accuracy, the inclusion of general relativistic corrections in orbit determination models becomes indispensable [12, 13, 14, 15, 16]. These corrections affect multiple aspects of orbit determination, including light-time equations, clock synchronization, station coordinates, and signal propagation delays. For instance, relativistic time delay effects, such as Shapiro delay, can introduce discrepancies of up to tens of microseconds per day, especially in deep-space navigation where signal propagation spans millions of kilometers [9]. Without these corrections, accumulated errors could lead to significant positional offsets, potentially corresponding to meter‐level range errors if uncorrected [9]. A fully consistent treatment across GCRS and BCRS is therefore essential not only to guarantee frame equivalence but also to extend GNSS capabilities to cislunar navigation.
In this work, we report on the implementation of the BCRS framework in JPL’s GipsyX [17], demonstrating millimeter-level closure results. We also derive the necessary coordinate and time transformations needed to define the Lunicentric Celestial Reference System (LCRS) for future Earth–Moon mission applications.
[…]
II Relativistic framework for processing GNSS data
II.1 Coordinate transformations between GCRS and BCRS
II.1.1 Direct transformations: GCRS $\rightarrow$ BCRS
According to the IAU conventions, the GCRS is realized by the geocentric metric tensor $G_{mn}$ in coordinates $(T,\mathbf{X})$, where $T\equiv{\tt TCG}$ denotes Geocentric Coordinate Time [18]. Following the approach of [9, 19], we express $G_{mn}$ to the order required for modern GNSS timing applications. Since current GNSS clocks reach fractional frequency stabilities of order $3\times 10^{-15}$, the recommended transformations [^1] The notational conventions employed in this paper are those used in [20, 1]. Letters from the second half of the Latin alphabet, m,n,=03m,n,...=0...3 denote spacetime indices. Greek letters $\alpha,\beta,...=1...3$ denote spatial indices. The metric $\gamma_{mn}$ is that of Minkowski spacetime with $\gamma_{mn}={\rm diag}(+1,-1,-1,-1)$ in the Cartesian representation. We employ the Einstein summation convention with indices being lowered or raised using $\gamma_{mn}$. We use powers of $G$ and negative powers of $c$ as bookkeeping devices for order terms. from the GCRS $(T={\tt TCG},\,\mathbf{X})$ to the BCRS $\bigl(t={\tt TCB},\,\mathbf{r}_{{\tt E}}=\mathbf{x}-\mathbf{x}_{{\tt E}}(t)\bigr)$ for GNSS data processing then take the form:
$T$ $=$ $t-c^{-2}\Big\{\int^{t}_{t_{0}}\Big({\textstyle\frac{1}{2}}v_{\tt E}^{2}+\sum_{{\tt B\not=E}}{GM_{\tt B}\over r_{\tt BE}}\Big)dt+(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{r}}_{\tt E})\Big\}+{\cal O}\Big(1.10\times 10^{-16}(t-t_{0})\Big),$ (1)
$\boldsymbol{\mathrm{X}}$ $=$ $\boldsymbol{\mathrm{r}}_{\tt E}+c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{r}}_{\tt E})\boldsymbol{\mathrm{v}}_{\tt E}+\sum_{{\tt B\not=E}}{GM_{\tt B}\over r_{\tt BE}}\boldsymbol{\mathrm{r}}_{\tt E}+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{r}}_{\tt E})\boldsymbol{\mathrm{r}}_{\tt E}-{\textstyle\frac{1}{2}}r^{2}_{\tt E}\boldsymbol{\mathrm{a}}_{\tt E}\Big\}+{\cal O}\Big(1.28\times 10^{-12}~{\rm m}\Big),$ (2)
where $\boldsymbol{\mathrm{x}}_{\tt E}$ and $\boldsymbol{\mathrm{v}}_{\tt E}=d\boldsymbol{\mathrm{x}}_{\tt E}/dt$ being the Earth’s position and velocity vectors in the BCRS, with their magnitudes known to be $x_{\tt E}=|\boldsymbol{\mathrm{x}}_{\tt E}|\equiv{\rm AU}\simeq 1.49\,597\,870\,700\times 10^{11}$ m, and $v_{\tt E}=|\boldsymbol{\mathrm{v}}_{\tt E}|=29\,784.7\,$ m/s, see Table 2. Note that in their most accurate form (see [18]), the time transformations (1) also include terms $\propto{\cal O}(c^{-4})$. The uncertainty in (1) comes from the omitted ${\cal O}(c^{-4})$ terms behaving as $\sim c^{-4}(-\frac{1}{8}v_{\tt E}^{2}-\frac{3}{2}v^{2}_{\tt E}GM_{\tt S}/r_{\tt E}+{\textstyle\frac{1}{2}}(GM_{\tt S}/r_{\tt E})^{2})\lesssim-1.10\times 10^{-16}=-9.50$ ps/d.
[…]
The acceleration-dependent terms in (2) may contribute up to $2.68\times 10^{-6}$ m at a ground station and $4.66\times 10^{-5}$ m at $r_{\tt GPS}$. Although these offsets are small, they become significant when comparing spacecraft accelerations in the BCRS and the GCRS. The dominant uncertainty in (2) stems from the neglected solar quadrupole moment, $J_{2}=2.25\times 10^{-7}$ [21, 22], which induces corrections of order $c^{-2}w_{2,{\tt S}}(t,{\bf x})\boldsymbol{\mathrm{r}}_{\tt E}\simeq c^{-2}(GM_{\tt S}J_{2}R_{\tt S}^{2}/{\rm AU}^{3})R_{\tt GPS}\sim 1.28\times 10^{-12}$ m, a value that is entirely negligible and therefore serves as a conservative error bound.
II.1.2 Inverse transformations: BCRS $\rightarrow$ GCRS
To compare the relativistic description of spacecraft acceleration across different coordinate reference systems, we also require the inverse transformations. Accordingly, the transformations from the BCRS back to the GCRS, inverse to (1)–(2), is obtained by noting that the relativistic time transformation at the GCRS origin ($\mathbf{X}=0$) is given as
$T=t^{*}-{1\over c^{2}}\int_{t_{0}}^{t^{*}}\Big({\textstyle{\textstyle{\textstyle{1\over 2}}}}v_{\tt E}^{2}+\sum_{{\tt B\not=E}}{GM_{\tt B}\over r_{\tt BE}}\Big)dt+{\cal O}\Big(1.10\times 10^{-16}(t^{*}-t_{0})\Big).$ (3)
and the TCG representation of the motion of the Earth with respect to the BCRS given as below
$\bar{\boldsymbol{\mathrm{x}}}_{\tt E}(T)=\boldsymbol{\mathrm{x}}_{\tt E}(t^{*}),$ (4)
These relations can be obtained either numerically or analytically. In the case of the direct transformations, (1)–(2), one must evaluate the integral in (1) by numerical quadrature or, where possible, in closed form. For the inverse transformations, the time transformations $t^{*}=t^{*}(T)$ is found by inverting (3), again using either numerical or analytical methods. Here, (3) follows directly from applying the GCRS time transformation, (1), at the geocenter ($\mathbf{X}=0$).
We can connect the Earth’s center world line, $\boldsymbol{\mathrm{x}}_{\tt E}(t)$, to an observer’s world line, $\boldsymbol{\mathrm{x}}(t)$, using a space-time curve that is orthogonal to both world lines. Note that in data processing programs, the relevant coordinate differences are not simply $\Delta\boldsymbol{\mathrm{r}}$, but rather the differences $\Delta\boldsymbol{\mathrm{r}}=\boldsymbol{\mathrm{x}}(t^{*})-\bar{\boldsymbol{\mathrm{x}}}_{\tt E}(t^{*})$, which are the coordinates at a fixed coordinate time, i.e., simultaneous in solar system barycenter coordinates.
Accordingly, in the inverse transformations of a geocentric position vector $\boldsymbol{\mathrm{X}}$ from the GCRS (compatible with $\tt TCG$, $\tt TAI$, or $\tt TT$ for ground stations) to the solar system barycentric frame (compatible with $\tt TDB$), one must account for the Lorentz contraction of $\boldsymbol{\mathrm{X}}$ and make relativistic adjustments to the scale of $\boldsymbol{\mathrm{X}}$, as illustrated below:
$t$ $=$ $T+{1\over c^{2}}\Big\{\int_{T_{0}}^{T}\Big({\textstyle{\textstyle{\textstyle{1\over 2}}}}v_{\tt E}^{2}+\sum_{{\tt B\not=E}}{GM_{\tt B}\over r_{\tt BE}}\Big)dT+({\boldsymbol{\mathrm{v}}}_{\tt E}\cdot{\boldsymbol{\mathrm{X}}})\Big\}+{\cal O}\Big(1.10\times 10^{-16}\Big)(T-T_{0}),$ (5)
$\boldsymbol{\mathrm{r}}_{\tt E}$ $=$ $\boldsymbol{\mathrm{X}}-c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\boldsymbol{\mathrm{v}}_{\tt E}+\sum_{{\tt B\not=E}}{GM_{\tt B}\over r_{\tt BE}}\boldsymbol{\mathrm{X}}+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\boldsymbol{\mathrm{X}}-{\textstyle\frac{1}{2}}X^{2}\boldsymbol{\mathrm{a}}_{\tt E}\Big\}+{\cal O}\Big(1.28\times 10^{-12}~{\rm m}\Big).$ (6)
As a result, we now have both direct (1)–(2) and inverse (5)–(6) coordinate transformations between BCRS and GCRS, which are provided with accuracy sufficient for processing GNSS data from either frame. Note that a similar set of transformations, albeit without acceleration-dependent terms, has been implemented for processing LLR data, as discussed in [23]. Special attention must be given when dealing with either ground stations or Earth-orbiting spacecraft, as the coordinates $\boldsymbol{\mathrm{X}}$ will differ [19]. Specifically, terms proportional to $1/c^{4}$ in the time transformations (1) and (5) are necessary if a frequency transfer accuracy of approximately $1.0\times 10^{-16}$ is required, and acceleration-dependent terms in position transformations (2) and (6) are needed for positional accuracy within a few millimeters.
II.2 Times $\tt TT$ and $\tt TDB$ and related scaling constants
II.2.1 Relativistic time scales
To compare clocks following different world-lines in the solar system one needs to account for the differences in relevant dynamical environment. We first consider the relationship between $\tt TT$ and $\tt TCG$. $\tt TT$ was defined by IAU Resolution A4 (1991) [24] as: ‘a time scale differing from $\tt TCG$ by a constant rate, with the unit of measurement of $\tt TT$ chosen so that it matches the SI second on the geoid.’ According to the transformation between proper and coordinate time, this constant rate is expressed as
$\frac{d{\tt TT}}{d{\tt TCG}}=1-\frac{1}{c^{2}}W_{\tt g}=1-L_{\tt G},$ (7)
where $W_{\tt g}$ is the combined gravitational and rotational potential on the geoid, determined as $W_{\tt g}=(62636856.0\pm 0.5)~{\rm m}^{2}{\rm s}^{-2}$ [25]. The IAU value for $L_{\tt G}$ is 6.969 $69\,290\,134\times 10^{-10}\approx 60.2$ microseconds/day $(\mu{\rm s/d})$, a defining constant as set by IAU 2000 Resolution B1.9, Table 1.1 in [5].
According to the IAU Resolutions [26], instead of coordinate time ${\tt T}={\tt TCG}$, spatial coordinates $\boldsymbol{\mathrm{X}}$ and mass factors $(GM)_{\tt TCG}$ related to GCRS, the following scaling of these quantiles is used [9, 19]
${\tt TT}={\tt TCG}-L_{\tt G}({\tt TCG}-{\tt T}_{0}),\qquad\boldsymbol{\mathrm{X}}_{\tt TT}=(1-L_{\tt G})\boldsymbol{\mathrm{X}}_{\tt TCG},\qquad(GM)_{\tt TT}=(1-L_{\tt G})(GM)_{\tt TCG}.$ (8)
Another constant $L_{\tt C}$ removes the constant rate between $\tt TCG$} and $\tt TCB$}. It is determined as the long time average of the rate computed from the full post-Newtonian version of (5) accurate to ${\cal O}(c^{-4})$ (see [9]):
$\Big<\frac{d{\tt TCG}}{d{\tt TCB}}\Big>=1-L_{\tt C}.$ (9)
According to the IAU 2000 Resolution B1.5, $L_{\tt C}=1.480\,826\,867\,41\times 10^{-8}\pm 2\times 10^{-17}\approx 1.279\,434\,4~\text{ms/d}\pm 1.7~\text{ps/d}$, introducing it as a fundamental constant (refer to Table 1.1 in [5]).
The IAU 2000 Resolution B1.5, recommended the following relationship between the constants $L_{\tt B}$, $L_{\tt G}$, and $L_{\tt C}$:
$\Big<\frac{d{\tt TT}}{d{\tt TCB}}\Big>=\Big(\frac{d{\tt TT}}{d{\tt TCG}}\Big)\Big<\frac{d{\tt TCG}}{d{\tt TCB}}\Big>\qquad\Rightarrow\qquad 1-L_{\tt B}=(1-L_{\tt G})(1-L_{\tt C}),$ (10)
where $L_{\tt B}$ is determined as $L_{\tt B}=L_{\tt G}+L_{\tt C}-L_{\tt G}L_{\tt C}=1.550\,519\,767\,72\times 10^{-8}\approx 1.34$ ms/d.
Finally, TDB is a timescale rescaled from TCB, as defined by IAU 2006 Resolution B3 and IAU 2009 Resolution 3 [27, 5], given by the following set of expressions:
${\tt TDB}={\tt TCB}-L_{\tt B}({\tt TCB}-{\tt T}_{0})+{\tt TDB}_{0},\qquad\boldsymbol{\mathrm{x}}_{\tt TDB}=(1-L_{\tt B})\boldsymbol{\mathrm{x}}_{\tt TCB},\qquad(GM)_{\tt TDB}=(1-L_{\tt B})(GM)_{\tt TCB},$ (11)
where, the defining constants $L_{\tt B}=1.550519768\times 10^{-8},{\tt T}_{0}=2443144.5003725\,\mathrm{JD}$, and ${\tt TDB}_{0}=-65.5\,\mu\mathrm{s},$ match those used in the JPL DE405 ephemeris [28]. This ensures that $\tt TDB$ advances at the same rate as $\tt TT$ at the geocenter. The offset ${\tt TDB}_{0}$ is chosen to align with the standard $(\tt TDB-TT)$ relation [29], which implies that $\tt TDB$} is not synchronized with $\tt TT$}, $\tt TCG$}, or $\tt TCB$} at 1977‑01‑01 00:00:32.184 TAI, at the geocenter (see discussion in [9, 19]).
Table 1: Reference time-scale constants and implied mean rates (IAU/IERS defining constants used in this work).
Symbol Definition Value (s/s) Mean rate per day Note
$L_{\tt G}$ $d{\tt TT}/d{\tt TCG}=1-L_{\tt G}$ 6.969 $69\,290\,134\times 10^{-10}$ 60.2146 $\mu{\rm s\,d}^{-1}$ IAU 2000 B1.9
$L_{\tt C}$ $\langle d{\tt TCG}/d{\tt TCB}\rangle=1-L_{\tt C}$ $1.480\,826\,867\,41\times 10^{-8}$ $1.279\,434\text{\,}\mathrm{ms}\text{\,}{\mathrm{d}}^{-1}$ IAU 2000 B1.5
$L_{\tt B}$ $1-L_{\tt B}=(1-L_{\tt G})(1-L_{\tt C})$ $1.550\,519\,767\,72\times 10^{-8}$ $1.339\,649\text{\,}\mathrm{ms}\text{\,}{\mathrm{d}}^{-1}$ IAU 2006/2009
$L_{\tt L}$ $d{\tt TL}/d{\tt TCL}=1-L_{\tt L}$ $3.1390541\times 10^{-11}$ 2.712143 $\mu{\rm s\,d}^{-1}$ this work, Eq. (88)
$L_{\tt H}$ $\langle d{\tt TCL}/d{\tt TCB}\rangle=1-L_{\tt H}$ $1.4825362\times 10^{-8}$ $1.280\,911\text{\,}\mathrm{ms}\text{\,}{\mathrm{d}}^{-1}$ this work, Eq. (91)
$L_{\tt M}$ $\langle d{\tt TL}/d{\tt TCB}\rangle=1-L_{\tt M}$ $1.485675294\times 10^{-8}$ $1.283\,624\text{\,}\mathrm{ms}\text{\,}{\mathrm{d}}^{-1}$ this work, Eq. (93)
II.2.2 Transformations from TT- to TDB-compatible quantities
To process the GNSS data, one needs an expression for the time transformations between GCRS ($\tt TT$, $\boldsymbol{\mathrm{X}}_{\tt TT}$) and BCRS ($\tt TDB$, $(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}=(\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}_{\tt E}(t))_{\tt TDB}$). By applying the relevant scaling relationships specified in (8) and (11), and introducing $(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}=(\boldsymbol{\mathrm{v}}_{\tt E}\cdot(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB})$, we have relationships in terms of the TDB-compatible quantities (see details in [19])
${\tt TT}$ $=$ ${\tt TDB}-{\tt TDB}_{0}+L_{\tt C}\Big({\tt TDB}-{\tt T}_{0}-{\tt TDB}_{0}\Big)$- (12)
- $c^{-2}\Big\{\int^{\tt TDB}_{{\tt T}_{0}+{\tt TDB}_{0}}\Big({\textstyle\frac{1}{2}}v_{\tt E}^{2}+\sum_{{\tt B\not=E}}{GM_{\tt B}\over r_{\tt BE}}\Big)dt+\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{r}}_{\tt E}\big)\Big\}_{\tt TDB}+{\cal O}\big(2.19\times 10^{-16}\big)\Big({\tt TDB}-{\tt T}_{0}-{\tt TDB}_{0}\Big),$~~
$\boldsymbol{\mathrm{X}}_{\tt TT}$ $=$ $\big(1+L_{\tt C}\big)\boldsymbol{\mathrm{r}}_{\tt E}{}_{\tt TDB}+c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{r}}_{\tt E})\boldsymbol{\mathrm{v}}_{\tt E}+\sum_{{\tt B\not=E}}{GM_{\tt B}\over r_{\tt BE}}\boldsymbol{\mathrm{r}}_{\tt E}+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{r}}_{\tt E})\boldsymbol{\mathrm{r}}_{\tt E}-{\textstyle\frac{1}{2}}r^{2}_{\tt E}\boldsymbol{\mathrm{a}}_{\tt E}\Big\}_{\tt TDB}+{\cal O}\Big(5.83\times 10^{-9}~{\rm m}\Big).$~~~~ (13)
Note that the errors in (12) are determined by the omitted products of the $L_{\tt C}^{2}$ term. If included, the error would be $\sim{\cal O}(1.46\times 10^{-16})$, resulting from the product of $L_{\tt C}$ and the external gravitational potential. Similarly, the error in (13) is set by the omitted term $L_{\tt C}^{2}(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}\simeq{\cal O}(5.83\times 10^{-9}~{\rm m})$.
II.2.3 Transformations from TDB- to TT-compatible quantities
Similarly, we can develop inverse transformations for transitioning from TDB-compatible to TT-compatible quantities. For this, we use expressions for the transforming from BCRS to GCRS given by (5)–(6). Then, applying the scaling relationships (8) and (11), we have:
${\tt TDB}$ $=$ $(1-L_{\tt C}){\tt TT}+L_{\tt C}{\tt T}_{0}+{\tt TDB}_{0}+{1\over c^{2}}\Big\{\int_{{\tt T}_{0}}^{\tt TT}\Big({\textstyle{\textstyle{\textstyle{1\over 2}}}}v_{\tt E}^{2}+\sum_{{\tt B\not=E}}{GM_{\tt B}\over r_{\tt BE}}\Big)dt+({\boldsymbol{\mathrm{v}}}_{\tt E}\cdot{\boldsymbol{\mathrm{X}}})\Big\}_{\tt TT}+{\cal O}\Big(1.46\times 10^{-16}\Big)\big({\tt TT}-{\tt T}_{0}\big),$~~~~~ (14)
$(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}$ $=$ $\big(1-L_{\tt C}\big)\boldsymbol{\mathrm{X}}_{\tt TT}-c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\boldsymbol{\mathrm{v}}_{\tt E}+\sum_{{\tt B\not=E}}{GM_{\tt B}\over r_{\tt BE}}\boldsymbol{\mathrm{X}}+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\boldsymbol{\mathrm{X}}-{\textstyle\frac{1}{2}}X^{2}\boldsymbol{\mathrm{a}}_{\tt E}\Big\}_{\tt TT}+{\cal O}\Big(5.83\times 10^{-9}~{\rm m}\Big).$~~~~~ (15)
Note that the error in (14) is set by the product of $L_{\tt C}$ and the external gravitational potential. Similar to (13), the errors in (15) are due to omitted $L_{\tt C}^{2}\boldsymbol{\mathrm{X}}_{\tt TT}\simeq 5.83\times 10^{-9}~{\rm m}$ term.
Result (14) yields the following useful derivative:
$\frac{d{\tt TDB}}{d{\tt TT}}$ $=$ $1-L_{\tt C}+c^{-2}\Big\{{\textstyle\frac{1}{2}}v_{\tt E}^{2}+\sum_{{\tt B\not=E}}{GM_{\tt B}\over r_{\tt BE}}+\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)+\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\Big\}_{\tt TT}+{\cal O}\Big(1.46\times 10^{-16}\Big).$ (16)
Now we have all the expressions needed to establish the transformation rules for velocities and accelerations.
II.3 Transforming positions, velocity and acceleration
II.3.1 Preliminary derivations
We now present the transformation required for the $\tt TT$-compatible GCRS position vector $\boldsymbol{\mathrm{X}}_{\tt TT}$, velocity $\boldsymbol{\mathrm{V}}_{\tt TT}=\dot{\boldsymbol{\mathrm{X}}}_{\tt TT}=d\boldsymbol{\mathrm{X}}/d{\tt TT}$, and acceleration $\boldsymbol{\mathrm{A}}_{\tt TT}=\ddot{\boldsymbol{\mathrm{X}}}_{\tt TT}=d\dot{\boldsymbol{\mathrm{X}}}_{\tt TT}/d{\tt TT}=d^{2}\boldsymbol{\mathrm{X}}/d{\tt TT}^{2}$. These transformations will convert these measurements to the BCRS with $\tt TDB$}-compatible position $(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}=(\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}_{\tt E})_{\tt TDB}$, velocity $(\dot{\boldsymbol{\mathrm{r}}}_{\tt E})_{\tt TDB}=d(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}/d{\tt TDB}=(\boldsymbol{\mathrm{v}}-\boldsymbol{\mathrm{v}}_{\tt E})_{\tt TDB}$, and acceleration $(\ddot{\boldsymbol{\mathrm{r}}}_{\tt E})_{\tt TDB}=d(\dot{\boldsymbol{\mathrm{r}}}_{\tt E})_{\tt TDB}/d{\tt TDB}=d^{2}(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}/d{\tt TDB}^{2}=(\boldsymbol{\mathrm{a}}-\boldsymbol{\mathrm{a}}_{\tt E})_{\tt TDB}$.
We begin with the position transformation that is given by (15), expressing it, for convenience, as below
$(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}$ $=$ $(\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}_{\tt E})_{\tt TDB}=\boldsymbol{\mathrm{X}}_{\tt TT}-\Delta\boldsymbol{\mathrm{X}}_{\tt TT},$ (17)
where the corrective term $\Delta\boldsymbol{\mathrm{X}}_{\tt TT}$ is of the order of ${\cal O}(c^{-2})$ and is given as
$\Delta\boldsymbol{\mathrm{X}}_{\tt TT}$ $=$ $L_{\tt C}\boldsymbol{\mathrm{X}}_{\tt TT}+c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\boldsymbol{\mathrm{v}}_{\tt E}+\gamma U_{\rm ext}\,\boldsymbol{\mathrm{X}}+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\boldsymbol{\mathrm{X}}-{\textstyle\frac{1}{2}}X^{2}\boldsymbol{\mathrm{a}}_{\tt E}\Big\}_{\tt TT}+{\cal O}\Big(5.83\times 10^{-9}~{\rm m}\Big),$ (18)
where, for convenience, we use
$U_{\rm ext}\equiv\sum_{{\tt B\not=E}}{GM_{\tt B}\over r_{\tt BE}}+{\cal O}(c^{-2}).$ (19)
Note that in (18) we also brought in the PPN parameter $\gamma$ as in [1]. We evaluate derivatives along $\tt TT$} unless noted: $\dot{U}_{\rm ext}=-\sum_{\tt B\not=E}GM_{\tt B}\,[(\boldsymbol{\mathrm{v}}_{\tt B}-\boldsymbol{\mathrm{v}}_{\tt E})\cdot(\boldsymbol{\mathrm{x}}_{\tt B}-\boldsymbol{\mathrm{x}}_{\tt E})]/r_{\tt BE}^{3}$ and $\ddot{U}_{\rm ext}$ analogously by differentiating once more; see (27)–(28).
Similarly, recognizing from (15) that $\boldsymbol{\mathrm{r}}_{\tt E}{}_{\tt TDB}=\boldsymbol{\mathrm{X}}_{\tt TT}+{\cal O}(c^{-2})$ and $\dot{\boldsymbol{\mathrm{r}}}_{\tt E}{}_{\tt TDB}=\dot{\boldsymbol{\mathrm{X}}}_{\tt TT}+{\cal O}(c^{-2})$ we present
$\frac{d{\tt TT}}{d{\tt TDB}}$ $=$ $1+L_{\tt C}-c^{-2}\Big\{{\textstyle\frac{1}{2}}v_{\tt E}^{2}+U_{\rm ext}+\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)+\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\Big\}_{\tt TT}+{\cal O}\Big(2.19\times 10^{-16}\Big).$ (20)
We can now develop the time derivatives of $(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}$:
$(\dot{\boldsymbol{\mathrm{r}}}_{\tt E})_{\tt TDB}$ $=$ $\frac{d(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}}{d{\tt TDB}}\equiv(\dot{\boldsymbol{\mathrm{x}}}-\boldsymbol{\mathrm{v}}_{\tt E})_{\tt TDB},\qquad(\ddot{\boldsymbol{\mathrm{r}}}_{\tt E})_{\tt TDB}=\frac{d^{2}(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}}{d{\tt TDB}^{2}}\equiv(\ddot{\boldsymbol{\mathrm{x}}}-\boldsymbol{\mathrm{a}}_{\tt E})_{\tt TDB}.$ (21)
For that, similarly to [30] and using expressions (17)–(18), we have
$\frac{d(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}}{d{\tt TDB}}$ $=$ $\frac{d\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}}\frac{d{\tt TT}}{d{\tt TDB}}-\frac{d\boldsymbol{\mathrm{\Delta}}\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}}\frac{d{\tt TT}}{d{\tt TDB}},$ (22)
$\frac{d^{2}(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}}{d{\tt TDB}^{2}}$ $=$ $\frac{d^{2}\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}^{2}}\Big(\frac{d{\tt TT}}{d{\tt TDB}}\Big)^{2}+\frac{d\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}}\frac{d^{2}{\tt TT}}{d{\tt TDB}^{2}}-\frac{d^{2}\boldsymbol{\mathrm{\Delta}}\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}^{2}}\Big(\frac{d{\tt TT}}{d{\tt TDB}}\Big)^{2}-\frac{d\boldsymbol{\mathrm{\Delta}}\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}}\frac{d^{2}{\tt TT}}{d{\tt TDB}^{2}}.$ (23)
From (20), we see that ${d{\tt TT}}/{d{\tt TDB}}=1+{\cal O}(c^{-2})$. In addition, $\Delta\boldsymbol{\mathrm{X}}_{\tt TT}$ is of the order of ${\cal O}(c^{-2})$. Given this, we may simplify (22) and (23), as below
$\frac{d(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}}{d{\tt TDB}}$ $=$ $\frac{d\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}}\frac{d{\tt TT}}{d{\tt TDB}}-\frac{d\boldsymbol{\mathrm{\Delta}}\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}}+{\cal O}(c^{-4}),$ (24)
$\frac{d^{2}(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}}{d{\tt TDB}^{2}}$ $=$ $\frac{d^{2}\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}^{2}}\Big(\frac{d{\tt TT}}{d{\tt TDB}}\Big)^{2}+\frac{d\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}}\frac{d^{2}{\tt TT}}{d{\tt TDB}^{2}}-\frac{d^{2}\boldsymbol{\mathrm{\Delta}}\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}^{2}}+{\cal O}(c^{-4}).$ (25)
Using (20), we have
$\frac{d{\tt TT}}{d{\tt TDB}}$ $=$ $1+L_{\tt C}-c^{-2}\Big\{{\textstyle\frac{1}{2}}v_{\tt E}^{2}+U_{\rm ext}+\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)+\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\Big\}_{\tt TT}+{\cal O}\Big(2.19\times 10^{-16}\Big),$ (26)
$\Big(\frac{d{\tt TT}}{d{\tt TDB}}\Big)^{2}$ $=$ $1+2L_{\tt C}-c^{-2}\Big\{{v_{\tt E}^{2}}+2U_{\rm ext}+2\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)+2\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\Big\}_{\tt TT}+{\cal O}\Big(4.38\times 10^{-16}\Big),$ (27)
$\frac{d^{2}{\tt TT}}{d{\tt TDB}^{2}}$ $=$ $-c^{-2}\Big\{(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{a}}_{\tt E})+\dot{U}_{\rm ext}+\big(\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)+2\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)+\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}}\big)\Big\}_{\tt TT}+{\cal O}\Big(2.91\times 10^{-23}\,{{\rm s}}^{-1}\Big),$~~~ (28)
where the error in ${d^{2}{\tt TT}}/{d{\tt TDB}^{2}}$ is set by the magnitude of the omitted $c^{-4}$ terms (12), see [19].
In addition, from (18), we compute
$\frac{d\boldsymbol{\mathrm{\Delta}}\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}}$ $=$ $L_{\tt C}\dot{\boldsymbol{\mathrm{X}}}_{\tt TT}+c^{-2}\Big\{{\textstyle\frac{1}{2}}\Big((\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{v}}_{\tt E}+\Big({\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})-({\boldsymbol{\mathrm{X}}}\cdot\dot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{a}}_{\tt E}+\Big((\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+\gamma U_{\rm ext}\Big)\dot{\boldsymbol{\mathrm{X}}}+$ (29)
$\thinspace+\,\Big((\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})+\gamma\dot{U}_{\rm ext}\Big)\boldsymbol{\mathrm{X}}-{\textstyle\frac{1}{2}}X^{2}\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\Big\}_{\tt TT}+{\cal O}\Big(8.49\times 10^{-13}~{\rm m/s}\Big),$~~~
where the errors in (29) are set by the term $L_{\tt C}^{2}(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}=L_{\tt C}^{2}\,\boldsymbol{\mathrm{X}}_{\tt TT}+{\cal O}(c^{-2})\simeq 5.83\times 10^{-9}~{\rm m}$ omitted in (29), that yields $L_{\tt C}^{2}\,\dot{\boldsymbol{\mathrm{X}}}_{\tt TT}\simeq 8.49\times 10^{-13}~{\rm m/s}$ for a MEO orbiter. Also, $\dot{U}_{\rm ext}={dU_{\rm ext}}/{d{\tt TT}}+{\cal O}(c^{-2})$ has the form
$\dot{U}_{\rm ext}\equiv\frac{dU_{\rm ext}}{d{\tt TT}}=-\sum_{{\tt B\not=E}}\frac{GM_{\tt B}\big((\boldsymbol{\mathrm{v}}_{\tt B}-\boldsymbol{\mathrm{v}}_{\tt E})\cdot(\boldsymbol{\mathrm{x}}_{\tt B}-\boldsymbol{\mathrm{x}}_{\tt E})\big)}{r^{3}_{\tt BE}}+{\cal O}(c^{-2}),$ (30)
with $GM_{\tt B},\boldsymbol{\mathrm{r}}_{\tt B}$ and $\boldsymbol{\mathrm{v}}_{\tt B}$ being the mass parameter and barycentric position and velocity vectors of body $\tt B$, $\tt E$ is the index for Earth and $r_{\tt BE}$ is the geocentric range of body $\tt B$ and $r_{\tt BE}=|\boldsymbol{\mathrm{x}}_{\tt B}-\boldsymbol{\mathrm{x}}_{\tt E}|$.
Finally, we use (29) to compute
$\frac{d^{2}\boldsymbol{\mathrm{\Delta}}\boldsymbol{\mathrm{X}}_{\tt TT}}{d{\tt TT}^{2}}$ $=$ $L_{\tt C}\ddot{\boldsymbol{\mathrm{X}}}_{\tt TT}+c^{-2}\Big\{{\textstyle\frac{1}{2}}\Big((\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+2(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})+(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{v}}_{\tt E}+\Big((\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})-(\dot{\boldsymbol{\mathrm{X}}}\cdot\dot{\boldsymbol{\mathrm{X}}})-({\boldsymbol{\mathrm{X}}}\cdot\ddot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{a}}_{\tt E}+$ (31)
$+$ $\Big({\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})-2({\boldsymbol{\mathrm{X}}}\cdot\dot{\boldsymbol{\mathrm{X}}})\Big)\dot{\boldsymbol{\mathrm{a}}}_{\tt E}+\Big(\gamma\ddot{U}_{\rm ext}+(\ddot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+2(\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{X}}+$
$+$ $2\Big(\gamma\dot{U}_{\rm ext}+(\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})\Big)\dot{\boldsymbol{\mathrm{X}}}+\Big(\gamma U_{\rm ext}+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\Big)\ddot{\boldsymbol{\mathrm{X}}}-{\textstyle\frac{1}{2}}X^{2}\ddot{\boldsymbol{\mathrm{a}}}_{\tt E}\Big\}_{\tt TT}+{\cal O}\Big(1.24\times 10^{-16}~{\rm m/s}^{2}\Big),$~~~~~~
where the errors in (31) are set by the omitted term $L_{\tt C}^{2}(\boldsymbol{\mathrm{r}}_{\tt E})_{\tt TDB}=L_{\tt C}^{2}\,\boldsymbol{\mathrm{X}}_{\tt TT}+{\cal O}(c^{-2})\simeq 5.83\times 10^{-9}~{\rm m}$ omitted in (29), that yields $L_{\tt C}^{2}\,\ddot{\boldsymbol{\mathrm{X}}}_{\tt TT}\simeq 1.24\times 10^{-16}~{\rm m/s}^{2}$ for a MEO orbiter. In addition, $\ddot{U}_{\rm ext}={d^{2}U_{\rm ext}}/{d{\tt TT}^{2}}+{\cal O}(c^{-2})$ has the form
$\ddot{U}_{\rm ext}\equiv\frac{d^{2}U_{\rm ext}}{d{\tt TT}^{2}}$ $=$ $-\sum_{{\tt B\not=E}}\frac{GM_{\tt B}}{r^{3}_{\tt BE}}\Big((\boldsymbol{\mathrm{a}}_{\tt B}-\boldsymbol{\mathrm{a}}_{\tt E})\cdot(\boldsymbol{\mathrm{x}}_{\tt B}-\boldsymbol{\mathrm{x}}_{\tt E})+$ (32)
$+~(\boldsymbol{\mathrm{v}}_{\tt B}-\boldsymbol{\mathrm{v}}_{\tt E})\cdot(\boldsymbol{\mathrm{v}}_{\tt B}-\boldsymbol{\mathrm{v}}_{\tt E})-\frac{3}{r^{2}_{\tt BE}}\big((\boldsymbol{\mathrm{v}}_{\tt B}-\boldsymbol{\mathrm{v}}_{\tt E})\cdot(\boldsymbol{\mathrm{x}}_{\tt B}-\boldsymbol{\mathrm{x}}_{\tt E})\big)^{2}\Big)+{\cal O}(c^{-2}),$
with $GM_{\tt B},\boldsymbol{\mathrm{r}}_{\tt B}$ and $\boldsymbol{\mathrm{v}}_{\tt B}$ being the mass parameter and barycentric position and velocity vectors of body $\tt B$, $\tt E$ is the index for Earth and $r_{\tt BE}$ is the geocentric range of body $\tt B$ and $r_{\tt BE}=|\boldsymbol{\mathrm{x}}_{\tt E}-\boldsymbol{\mathrm{x}}_{\tt B}|$.
II.3.2 Transforming velocity and acceleration
Now, we can present the expressions that describe the velocity and acceleration transformations. To do this, starting with the velocity, we use (22), (24), (26), and (29), which yield
$(\dot{\boldsymbol{\mathrm{x}}}-\dot{\boldsymbol{\mathrm{x}}}_{\tt E})_{\tt TDB}$ $=$ $\dot{\boldsymbol{\mathrm{X}}}_{\tt TT}-c^{-2}\Big\{\Big({\textstyle\frac{1}{2}}v_{\tt E}^{2}+(1+\gamma)U_{\rm ext}+2\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)+\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\Big)\dot{\boldsymbol{\mathrm{X}}}+{\textstyle\frac{1}{2}}\Big((\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{v}}_{\tt E}+$ (33)
$\thinspace+~\Big({\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})-({\boldsymbol{\mathrm{X}}}\cdot\dot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{a}}_{\tt E}+\Big((\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})+\gamma\dot{U}_{\rm ext}\Big)\boldsymbol{\mathrm{X}}-{\textstyle\frac{1}{2}}X^{2}\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\Big\}_{\tt TT}+{\cal O}\Big(8.49\times 10^{-13}~{\rm m/s}\Big).$~~~~~
Note that all the quantities on the right-hand side of (33) are TT-compatible.
Similarly, we have the expression for transforming the acceleration between the two reference systems. For this, we will use (22), (25), (27)–(28), and (31), which yield the following:
$(\ddot{\boldsymbol{\mathrm{x}}}-\ddot{\boldsymbol{\mathrm{x}}}_{\tt E})_{\tt TDB}$ $=$ $\ddot{\boldsymbol{\mathrm{X}}}_{\tt TT}+L_{\tt C}\ddot{\boldsymbol{\mathrm{X}}}_{\tt TT}-c^{-2}\Big\{\Big({v_{\tt E}^{2}}+(2+\gamma)U_{\rm ext}+3\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)+2\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\Big)\ddot{\boldsymbol{\mathrm{X}}}+$ (34)
$+$ $\Big((\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{a}}_{\tt E})+(1+2\gamma)\dot{U}_{\rm ext}+3\big(\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)+4\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)+\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}}\big)\Big)\dot{\boldsymbol{\mathrm{X}}}+$
$+$ ${\textstyle\frac{1}{2}}\Big((\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+2(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})+(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{v}}_{\tt E}+\Big((\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})-(\dot{\boldsymbol{\mathrm{X}}}\cdot\dot{\boldsymbol{\mathrm{X}}})-({\boldsymbol{\mathrm{X}}}\cdot\ddot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{a}}_{\tt E}+$
$+$ $\Big({\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})-2({\boldsymbol{\mathrm{X}}}\cdot\dot{\boldsymbol{\mathrm{X}}})\Big)\dot{\boldsymbol{\mathrm{a}}}_{\tt E}+\Big(\gamma\ddot{U}_{\rm ext}+(\ddot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+2(\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{X}}-{\textstyle\frac{1}{2}}X^{2}\ddot{\boldsymbol{\mathrm{a}}}_{\tt E}\Big\}_{\tt TT}+$
$+$ ${\cal O}\Big(1.24\times 10^{-16}~{\rm m/s}^{2}\Big),$
where the error terms in these expressions are set by the omitted $L^{2}_{\tt C}$ terms.
II.4 Practical form for the relevant expressions
Results (17)–(18), (33), and (34) have many terms that may be too small. We estimate the magnitudes of the terms in these expressions using Table 2, which provides the nominal values for the quantities involved. This will help us understand the changes in station or satellite coordinates relative to the Earth’s center of mass in the barycentric frame. We also recognize that the solar gravitational potential provides the largest contribution to $U_{\rm ext}$.
Table 2: Approximate values for various parameters involved.
$\gamma$ Parametrized post-Newtonian (PPN) parameter 1
$GM_{\tt S}$ Solar gravitational constant $1.32712442099\times 10^{20}$ ${\rm m}^{3}{\rm s}^{-2}$
$r_{\tt E}$ Earth’s position relative to barycenter $1\,{\rm AU}=149\,597\,870\,700.0~\mathrm{m}$
${\boldsymbol{\mathrm{v}}}_{\tt E}$ Earth’s velocity relative to barycenter $29.78\times 10^{3}$ m/s
Ground-based station:
$\boldsymbol{\mathrm{X}}_{\tt S}$ position relative to the center of the Earth $6378.1\times 10^{3}$ m
$\dot{\boldsymbol{\mathrm{X}}}_{\tt S}$ velocity relative to the center of the Earth $({2\pi}/{\rm day})R_{\tt E}=463.83$ m/s
$\ddot{\boldsymbol{\mathrm{X}}}_{\tt S}$ acceleration relative to the center of the Earth   $({2\pi}/{\rm day})^{2}R_{\tt E}=3.37\times 10^{-2}$ m/s2
MEO spacecraft at altitude of $h=20200$ km
$\boldsymbol{\mathrm{X}}_{\tt MEO}$ position relative to the center of the Earth $26578.1\times 10^{3}$ m
$\dot{\boldsymbol{\mathrm{X}}}_{\tt MEO}$ velocity relative to the center of the Earth $3.87\times 10^{3}$ m/s
$\ddot{\boldsymbol{\mathrm{X}}}_{\tt MEO}$ acceleration relative to the center of the Earth $0.56$ m/s2
As the values for a MEO (GPS‑like) satellite are much larger than those for a ground-based station, we consider the transformations for a MEO orbiter (i.e., a GPS spacecraft), with the position, velocity and the acceleration present in (17)–(18), (33), and (34) take the relevant values of MEO orbiter (i.e., a GPS spacecraft), namely $({\boldsymbol{\mathrm{X}}},\dot{\boldsymbol{\mathrm{X}}},\ddot{\boldsymbol{\mathrm{X}}})\rightarrow({\boldsymbol{\mathrm{X}}}_{\tt MEO},\dot{\boldsymbol{\mathrm{X}}}_{\tt MEO},\ddot{\boldsymbol{\mathrm{X}}}_{\tt MEO})$, that are given in Table 2. With these values, we use (18) to estimate contribution of the various terms in the position transformation:
$\Big(L_{\tt C}+c^{-2}\gamma U_{\rm ext}\Big){\boldsymbol{\mathrm{X}}}\approx 0.66\,{\rm m},\qquad c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot{\boldsymbol{\mathrm{X}}})\boldsymbol{\mathrm{v}}_{\tt E}\Big\}\approx 0.13\,{\rm m},\qquad c^{-2}\Big\{(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\boldsymbol{\mathrm{X}}-{\textstyle\frac{1}{2}}X^{2}\boldsymbol{\mathrm{a}}_{\tt E}\Big\}\lesssim 7.00\times 10^{-5}~{\rm m}.$ (35)
Note that the acceleration-dependent terms in (35) are very small and are currently considered practically irrelevant. Consequently, a reasonable approximation of (17)–(18) for a MEO orbiter would be, as outlined in [1],
$(\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}_{\tt E})_{\tt TDB}$ $=$ $\boldsymbol{\mathrm{X}}_{\tt TT}-L_{\tt C}\boldsymbol{\mathrm{X}}_{\tt TT}-c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\boldsymbol{\mathrm{v}}_{\tt E}+\gamma U_{\rm ext}\,\boldsymbol{\mathrm{X}}\Big\}_{\tt TT}+{\cal O}\Big(7.01\times 10^{-5}~{\rm m}\Big),$ (36)
where the error bound is provided by the omitted acceleration-dependent terms shown in (35). Note that the same expression is valid for a ground-based station with the error term being of the order of ${\cal O}\big(4.03\times 10^{-6}~{\rm m}\big)$.
Equation (36) is referred to as the modified Lorentz transformation. It incorporates the special relativity Lorentz transformation due to Earth’s barycentric velocity, general relativity effects from the gravitational potential, and scaling factors. This transformation affects not only the length of position vectors but also alters the angle between position vectors and the Earth’s velocity vector (aberration). This equation includes more corrective terms than those in ([1], Eq. (4-10)), which JPL’s interplanetary orbit determination software uses for converting station positions from TT-compatible quantities in GCRS to TDB-compatible quantities in BCRS when modeling observables.
Next, we evaluate the terms present in the velocity transformation in (33) as below
$c^{-2}\Big\{\Big({\textstyle\frac{1}{2}}v_{\tt E}^{2}+(1+\gamma)U_{\rm ext}\Big)\dot{\boldsymbol{\mathrm{X}}}\Big\}\approx 9.56\times 10^{-5}\,{\rm m/s},\qquad~~~c^{-2}\Big\{\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\dot{\boldsymbol{\mathrm{X}}}\Big\}$ $\approx$ $4.96\times 10^{-6}\,{\rm m/s},$ (37)
$c^{-2}\Big\{2\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)\dot{\boldsymbol{\mathrm{X}}}\Big\}\approx 1.36\times 10^{-8}\,{\rm m/s},\qquad c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})\boldsymbol{\mathrm{v}}_{\tt E}\Big\}$ $\approx$ $3.82\times 10^{-5}\,{\rm m/s}$~~~ (38)
$c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\boldsymbol{\mathrm{v}}_{\tt E}\Big\}\approx 2.61\times 10^{-8}\,{\rm m/s},\qquad~~~~c^{-2}\Big\{({\boldsymbol{\mathrm{X}}}\cdot\dot{\boldsymbol{\mathrm{X}}})\boldsymbol{\mathrm{a}}_{\tt E}\Big\}$ $\lesssim$ $3.40\times 10^{-9}\,{\rm m/s},$ (39)
$c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\boldsymbol{\mathrm{a}}_{\tt E}\Big\}\approx 2.61\times 10^{-8}\,{\rm m/s},\qquad\hskip 25.0ptc^{-2}\Big\{\gamma\dot{U}_{\rm ext}\boldsymbol{\mathrm{X}}\Big\}$ $\approx$ $5.22\times 10^{-8}\,{\rm m/s},$ (40)
$c^{-2}\Big\{\Big((\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{X}}\Big\}\approx 6.81\times 10^{-9}\,{\rm m/s},\qquad~~~~~~~\,c^{-2}\Big\{{\textstyle\frac{1}{2}}X^{2}\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\Big\}$ $\approx$ $9.28\times 10^{-12}\,{\rm m/s}.$ (41)
As a result, limiting the terms at $4.96\times 10^{-6}$ m/s, from (33), we have the following for the velocity transformation:
$(\dot{\boldsymbol{\mathrm{x}}}-\dot{\boldsymbol{\mathrm{x}}}_{\tt E})_{\tt TDB}$ $=$ $\dot{\boldsymbol{\mathrm{X}}}_{\tt TT}-c^{-2}\Big\{\Big({\textstyle\frac{1}{2}}v_{\tt E}^{2}+(1+\gamma)U_{\rm ext}+\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\Big)\dot{\boldsymbol{\mathrm{X}}}+{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})\boldsymbol{\mathrm{v}}_{\tt E}\Big\}_{\tt TT}+{\cal O}\Big(1.29\times 10^{-7}~{\rm m/s}\Big),$ (42)
where the error bound results from the omitted acceleration-dependent and $\dot{U}_{\rm ext}$ terms. It is worth noting that result (42) includes two additional terms compared to [1]. The same expression is also applicable to ground-based stations, with the corresponding error term being on the order of ${\cal O}(9.73\times 10^{-8}~{\rm m/s})$.
Finally, we evaluate the terms involved in the acceleration transformations (34) as below
$L_{\tt C}\ddot{\boldsymbol{\mathrm{X}}}\approx 8.36\times 10^{-9}\,{\rm m/s}^{2},\qquad~~\,c^{-2}\Big\{{\textstyle\frac{1}{2}}\Big(2(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})+(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{v}}_{\tt E}\Big\}$ $\approx$ $2.97\times 10^{-9}\,{\rm m/s}^{2},$ (43)
$c^{-2}\Big\{\Big({v_{\tt E}^{2}}+(2+\gamma)U_{\rm ext}+3\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)+2\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\Big)\ddot{\boldsymbol{\mathrm{X}}}\Big\}$ $\approx$ $2.37\times 10^{-8}\,{\rm m/s}^{2},$ (44)
$c^{-2}\Big\{\Big((\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{a}}_{\tt E})+(1+2\gamma)\dot{U}_{\rm ext}+4\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)+\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}}\big)\Big)\dot{\boldsymbol{\mathrm{X}}}\Big\}$ $\approx$ $7.58\times 10^{-10}\,{\rm m/s}^{2},$ (45)
$c^{-2}\Big\{3\big(\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)\dot{\boldsymbol{\mathrm{X}}}\Big\}\approx 8.11\times 10^{-15}\,{\rm m/s}^{2},\qquad~~~c^{-2}\Big\{{\textstyle\frac{1}{2}}(\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})\boldsymbol{\mathrm{v}}_{\tt E}\Big\}$ $\approx$ $1.40\times 10^{-14}\,{\rm m/s}^{2},$ (46)
$c^{-2}\Big\{\Big((\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})-(\dot{\boldsymbol{\mathrm{X}}}\cdot\dot{\boldsymbol{\mathrm{X}}})-({\boldsymbol{\mathrm{X}}}\cdot\ddot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{a}}_{\tt E}\Big\}$ $\approx$ $9.60\times 10^{-12}\,{\rm m/s}^{2},$ (47)
$c^{-2}\Big\{\Big({\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})-2({\boldsymbol{\mathrm{X}}}\cdot\dot{\boldsymbol{\mathrm{X}}})\Big)\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\Big\}$ $\approx$ $1.58\times 10^{-14}\,{\rm m/s}^{2},$ (48)
$c^{-2}\Big\{\Big(\gamma\ddot{U}_{\rm ext}+(\ddot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+2(\dot{\boldsymbol{\mathrm{a}}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{X}}\Big\}$ $\approx$ $3.70\times 10^{-14}\,{\rm m/s}^{2},$ (49)
$c^{-2}\Big\{(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}})\boldsymbol{\mathrm{X}}\Big\}\approx 9.90\times 10^{-13}\,{\rm m/s}^{2},\qquad~~~~~~~~~c^{-2}\Big\{{\textstyle\frac{1}{2}}X^{2}\ddot{\boldsymbol{\mathrm{a}}}_{\tt E}\Big\}$ $\approx$ $7.35\times 10^{-18}\,{\rm m/s}^{2}.$ (50)
As a result, limiting the terms to ${\cal O}(1\times 10^{-13}\,{\rm m/s}^{2})$, we have the following expression to transform acceleration:
$(\ddot{\boldsymbol{\mathrm{x}}}-\ddot{\boldsymbol{\mathrm{x}}}_{\tt E})_{\tt TDB}$ $=$ $\ddot{\boldsymbol{\mathrm{X}}}_{\tt TT}+L_{\tt C}\ddot{\boldsymbol{\mathrm{X}}}_{\tt TT}-c^{-2}\Big\{\Big({v_{\tt E}^{2}}+(2+\gamma)U_{\rm ext}+3\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}}\big)+2\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\Big)\ddot{\boldsymbol{\mathrm{X}}}+$ (51)
$+$ $\Big((\boldsymbol{\mathrm{v}}_{\tt E}\cdot\boldsymbol{\mathrm{a}}_{\tt E})+(1+2\gamma)\dot{U}_{\rm ext}+4\big(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)+\big(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}}\big)\Big)\dot{\boldsymbol{\mathrm{X}}}+\Big((\boldsymbol{\mathrm{a}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})+{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{v}}_{\tt E}+$
$+$ $\Big((\boldsymbol{\mathrm{a}}_{\tt E}\cdot\boldsymbol{\mathrm{X}})+(\boldsymbol{\mathrm{v}}_{\tt E}\cdot\dot{\boldsymbol{\mathrm{X}}})-(\dot{\boldsymbol{\mathrm{X}}}\cdot\dot{\boldsymbol{\mathrm{X}}})-({\boldsymbol{\mathrm{X}}}\cdot\ddot{\boldsymbol{\mathrm{X}}})\Big)\boldsymbol{\mathrm{a}}_{\tt E}+(\boldsymbol{\mathrm{a}}_{\tt E}\cdot\ddot{\boldsymbol{\mathrm{X}}})\boldsymbol{\mathrm{X}}\Big\}_{\tt TT}+{\cal O}\Big(6.68\times 10^{-14}~{\rm m/s}^{2}\Big),$
where the error bound is from the group of the omitted terms involving $\dot{\boldsymbol{\mathrm{a}}}_{\tt E}$ and $\ddot{U}_{\rm ext}$ derivatives. It is worth noting that result (51) includes several additional terms compared to [1]. Also, the same expression is also applicable to ground-based stations, with the corresponding error term being on the order of ${\cal O}\big(5.27\times 10^{-15}~{\rm m/s}^{2}\big)$.
Therefore, considering realistic measurement accuracies, we refer to the magnitudes of various terms conducted for a MEO orbiter. Consequently, the expressions (36), (42), and (51) are recommended for position, velocity, and acceleration transformations when GNSS data processing is concerned. These results represent the set of the expressions recommended to transform position, velocity and acceleration vectors as needed to process GNSS data from either GCRS or BCRS. These formulae are accurate to order $c^{-2}$ and the corrective terms can be evaluated using for $\boldsymbol{\mathrm{X}}_{\tt TT},\boldsymbol{\mathrm{V}}_{\tt TT}$ and $\boldsymbol{\mathrm{A}}_{\tt TT}$ in TT-compatible GCRS or TDB-compatible BCRS quantities translated to the geocenter. These formulae are accurate to ${\cal O}(c^{-2})$. We adopt GR (i.e., $\gamma=1$) for all numerical results.
Recommended transformations: For cm/ps-class GNSS, use Eq. (36) for position with the two leading $c^{-2}$ terms (scale and Lorentz), Eq. (42) for velocity with the three leading $c^{-2}$ terms, and Eq. (51) for acceleration retaining terms $\gtrsim 10^{-13}\,\mathrm{m\,s^{-2}}$. These screened sets were used to produce the closures in Fig. 1.
The inverse transformations to (36), (42), and (51) are obtained by interchanging $({\boldsymbol{\mathrm{r}}}_{\tt E})_{\tt TDB}=(\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}_{\tt E})_{\tt TDB}$ with $\boldsymbol{\mathrm{X}}_{\tt TT}$, $(\dot{\boldsymbol{\mathrm{r}}}_{\tt E})_{\tt TDB}=(\dot{\boldsymbol{\mathrm{x}}}-\dot{\boldsymbol{\mathrm{x}}}_{\tt E})_{\tt TDB}=(\boldsymbol{\mathrm{v}}-\boldsymbol{\mathrm{v}}_{\tt E})_{\tt TDB}$ with $\boldsymbol{\mathrm{V}}_{\tt TT}$, and $(\ddot{\boldsymbol{\mathrm{r}}}_{\tt E})_{\tt TDB}=(\ddot{\boldsymbol{\mathrm{x}}}-\ddot{\boldsymbol{\mathrm{x}}}_{\tt E})_{\tt TDB}=(\boldsymbol{\mathrm{a}}-\boldsymbol{\mathrm{a}}_{\tt E})_{\tt TDB}$ with $\boldsymbol{\mathrm{A}}_{\tt TT}$ in all 3 equations and inverting the sign of all corrective terms, as evident from (13) and (15).
For 24 h spans with cm/ps-class observables, enforce
$\sup_{t}\|\delta\mathbf{r}\|\leq 7.01\times 10^{-5}\,\mathrm{m},\qquad\sup_{t}\|\delta\mathbf{v}\|\leq 1.29\times 10^{-7}\,\mathrm{m\,s^{-1}},\qquad\sup_{t}\|\delta\mathbf{a}\|\leq 6.68\times 10^{-14}\,\mathrm{m\,s^{-2}},$
matching the remainders from Eqs. (36), (42), and (51) used in our closures. Document these bounds with the same ephemerides to enable external regression.
Finally, mass parameters of celestial bodies, $GM$, must also be transformed between their value in the TT-compatible GCRS, $(GM)_{\tt TT}$, and the TDB-compatible value in BCRS, $(GM)_{\tt TDB}$, as:
$(GM)_{\tt TDB}=(1-L_{\tt C})(GM)_{\tt TT},$ (52)
which may be obtained from (8) and (11), by equating $(GM)_{\tt TCB}=(GM)_{\tt TCG}$ and using (10).
We stress that dynamical parameters tied to a TDB‑compatible ephemeris (e.g., DE440) must not be re‑scaled when used with Eqs. (51)–(52); only the labels of the time arguments change. The IAU constants used are $L_{\tt G}=6.969290134\times 10^{-10}$ and $L_{\tt B}=1.550519768\times 10^{-8}$, with the conventional ${\tt TDB}_{0}\simeq-65.5~\mu\mathrm{s}$.
Implementation note (GM scaling & ephemeris argument): This work adopts TDB as the ephemeris argument. For JPL DE430/DE440-family kernels, the delivered $(GM)$ values are already TDB-consistent; use Eq. (52) only when converting a TT-referenced constant, i.e., $(GM)_{\tt TDB}=(1-L_{\tt C})(GM)_{\tt TT}$ and inversely $(GM)_{\tt TT}=(1-L_{\tt C})^{-1}(GM)_{\tt TDB}$. Do not rescale any $(GM)$ that is tied to a TDB-based kernel.
II.5 Minimal cislunar case study: NRHO-like geometry
[…]
State propagation and transformations:
Propagate the spacecraft state in the BCRS with the same dynamical model used for the MEO test (relativistic point-mass terms and geopotential terms exactly as in the main text), and adopt the screened operational forms for position, velocity, and acceleration, Eqs. (36), (42), and (51), respectively. Map spacecraft and ground states to the LCRS using Eqs. (108) and (109) for position and velocity; accelerations follow the same screening logic as in the Earth case (see Table 3 for configuration details.)
Observables and light-time:
Compute one-way light-time between a ground station and the cislunar spacecraft with (64). Retain Sun and Earth Shapiro terms only; other bodies are negligible at the accuracy targeted here. (Note that most of the solar term behaves as a range scale and cancels in differenced observables, but keep it for absolute light-time consistency at the centimeter level over $\Delta T$.) Use the same TT/TDB handling and GM scaling guidance as in Sec. II.4 to avoid double-scaling when using TDB-consistent ephemerides.
[…]
Define the cislunar frame-closure for position and velocity over an analysis span $\Delta T$ as
$\Delta\boldsymbol{\mathrm{r}}(t)\equiv\boldsymbol{\mathrm{r}}_{\tt BCRS\to LCRS\to BCRS}(t)-\boldsymbol{\mathrm{r}}_{\tt BCRS}(t),\qquad\Delta\boldsymbol{\mathrm{v}}(t)\equiv\boldsymbol{\mathrm{v}}_{\tt BCRS\to LCRS\to BCRS}(t)-\boldsymbol{\mathrm{v}}_{\tt BCRS}(t),$ (53)
and require that the sup norms over $t\in[t_{0},t_{0}+\Delta T]$ do not exceed the screening remainders already specified under Eqs. (36) and (42), respectively. This ensures that all omitted terms remain below the stated operational error budgets when operating in cislunar geometries.
Table 3: Configuration for the cislunar check in Sec. II.5 (NRHO-like geometry, Earth relay).
Ephemerides and time scales DE440 (TDB); TT/TDB transformations as in Eqs. (12)–(15)
State propagation frame BCRS
Dynamics (spacecraft) Same relativistic and geopotential model as used for MEO closure
Operational transformations (Earth) Position/velocity/accel: Eqs. (36), (42), (51)
LCRS transformations (Moon) Position/velocity (108), (109); acceleration (110).
Light-time and delays Eq. (64); Shapiro bodies: Sun, Earth
Closure metric Eq. (53); thresholds tied to remainders of Eqs. (36), (42)
Span $\Delta T$ 24 h (nominal); longer spans permissible for operational testing
Practical notes:
(i) Use the same station coordinates and Earth-orientation handling as in the MEO test so that any differences trace solely to the frame transformations. (ii) When constructing differenced observables, the solar Shapiro contribution behaves dominantly as a scale; retain it for absolute range, document its cancellation otherwise. (iii) Publish the reference states and resulting $\Delta\boldsymbol{\mathrm{r}},\Delta\boldsymbol{\mathrm{v}}$ statistics for this case alongside those of the MEO closure to enable external regression tests.
II.6 Transformation between the proper time and TDB in BCRS
II.6.1 Direct transformations: TDB $\rightarrow$ proper time
Considering GNSS applications in cislunar environment we may use a truncated version of the metric tensor of the BCRS (see details in [19]). Then, to sufficient accuracy, the transformation between the proper time of an observer $\tau$ and the coordinate time in the BCRS, $\tt TCB$, is given as below:
$\frac{d\tau}{d{\tt TCB}}=1-{1\over c^{2}}\Big\{{\textstyle\frac{1}{2}}v_{0}^{2}+\sum_{\tt B}{GM_{\tt B}\over r_{\tt B0}}\Big(1-J_{\tt B2}\Big(\frac{R_{\tt B}}{r_{\tt B0}}\Big)^{2}P_{20}(\cos\theta_{\tt B})\Big)\Big\}_{\tt TCB}+{\cal O}\Big(8.77\times 10^{-17}\Big),$ (54)
where ${\boldsymbol{\mathrm{x}}}_{0}$, $\boldsymbol{\mathrm{v}}_{0}=\dot{\boldsymbol{\mathrm{x}}}_{0}$, and $v_{0}=|\boldsymbol{\mathrm{v}}_{0}|$ represent the BCRS position and velocity vectors of the spacecraft; $\boldsymbol{\mathrm{r}}_{\tt B0}=\boldsymbol{\mathrm{x}}_{B}-\boldsymbol{\mathrm{x}}_{0}$ is the separation vector between a body $B$ and the spacecraft, $r_{\tt B0}=|\boldsymbol{\mathrm{r}}_{\tt B0}|$; also we define $\cos\theta_{\tt B}=(\boldsymbol{\mathrm{n}}_{\tt B0}\cdot\boldsymbol{\mathrm{s}}_{\tt B})$, where $\boldsymbol{\mathrm{s}}_{\tt B}$ is the unit vector along the axis of rotation for the body B. Note that we kept only quadrupole term in $w_{\tt L}$, as contributions of other multipole harmonics of the gravitational fields are typically very small. Ultimately, the error bound in (54) came from the omitted contribution of the $1/c^{4}$–terms (see [19]), that for a GPS spacecraft evaluated to have magnitude of $c^{{}_{4}}(-{\textstyle{1\over 8}}v^{4}-{\textstyle{3\over 2}}v^{2}w_{0}+4(\boldsymbol{\mathrm{v}}\cdot\boldsymbol{\mathrm{w}})+{\textstyle{\textstyle{\textstyle{1\over 2}}}}w_{0}^{2}+\Delta)\simeq 8.77\times 10^{-17}.$ As a result, we may neglect these terms for GNSS applications in cislunar environment.
Taking into account (11), the relationships between the $\tt TDB$– and $\tt TCB$–compatible quantities yield a relation for the rates of an observer’s proper time and $\tt TDB$. Specifically, by applying (11) and using the chain rule for derivatives, $d\tau/d{\tt TDB}=(d\tau/d{\tt TCB})(d{\tt TCB}/d{\tt TDB})=(d\tau/d{\tt TCB})/(1-L_{\tt B})$, we obtain the following result:
$\frac{d\tau}{d{\tt TDB}}=\frac{1}{1-L_{\tt B}}\Big(1-{1\over c^{2}}\Big\{{\textstyle\frac{1}{2}}v_{0}^{2}+\sum_{\tt B}{GM_{\tt B}\over r_{\tt B0}}\Big(1-\frac{J_{\tt B2}R_{\tt B}^{2}}{2r_{\tt B0}^{2}}\big(3(\boldsymbol{\mathrm{n}}_{\tt B0}\cdot\boldsymbol{\mathrm{s}}_{\tt B})^{2}-1\big)\Big)\Big\}_{\tt TDB}\Big)+{\cal O}\Big(8.77\times 10^{-17}\Big).$ (55)
Integrating (55) provides an expression that relates the interval of the elapsed proper time $\tau$ during a particular interval of $\tt TDB\in[\tt TDB,\tt TDB_{\tau}]$:
$\tau$ $=$ $\frac{1}{1-L_{\tt B}}\Big({\tt TDB}-{\tt TDB}_{\tau}-\frac{1}{c^{2}}\int^{\tt TDB}_{{\tt TDB_{\tau}}}\Big\{{\textstyle\frac{1}{2}}v_{0}^{2}+\sum_{\tt B}{GM_{\tt B}\over r_{\tt B0}}\Big(1-\frac{J_{\tt B2}R_{\tt B}^{2}}{2r_{\tt B0}^{2}}\big(3(\boldsymbol{\mathrm{n}}_{\tt B0}\cdot\boldsymbol{\mathrm{s}}_{\tt B})^{2}-1\big)\Big)\Big\}_{\tt TDB}dt\Big)+$ (56)
$+$ ${\cal O}\Big(8.77\times 10^{-17}\Big)\big({\tt TDB}-{\tt TDB}_{\tau}\big),$
where $\boldsymbol{\mathrm{r}}_{0}$ and $\boldsymbol{\mathrm{v}}_{0}=d\boldsymbol{\mathrm{r}}_{0}/dt$ are the barycentric position and velocity of the spacecraft, $\boldsymbol{\mathrm{r}}_{\tt B0}=\boldsymbol{\mathrm{r}}_{0}-\boldsymbol{\mathrm{r}}_{\tt B}$ is its position relative to a body $B$, with $\boldsymbol{\mathrm{n}}_{\tt B0}=\boldsymbol{\mathrm{r}}_{\tt B0}/r_{\tt B0}$ being the unit vector in this direction, and $\boldsymbol{\mathrm{s}}_{\tt B}$ is the unit vector alone the axis of the rotation, $\tt TDB_{\tau}$ is the beginning of the integration interval.
II.6.2 Inverse transformations: proper time $\rightarrow$ TDB
Similarly to (54), we determine the inverse transformation from the proper time to the TCB:
$\frac{d{\tt TCB}}{d\tau}=1+{1\over c^{2}}\Big\{{\textstyle\frac{1}{2}}v^{2}_{0}+\sum_{\tt B}{GM_{\tt B}\over r_{\tt B0}}\Big(1-J_{\tt B2}\Big(\frac{R_{\tt B}}{r_{\tt B0}}\Big)^{2}P_{20}(\cos\theta_{\tt B})\Big)\Big\}_{\tt\tau}+{\cal O}\Big(1.37\times 10^{-16}\Big),$ (57)
where then error bound came from the omitted contribution of the $c^{-4}$-terms in the inverse transformation that were estimated to contribute effect on the order of $c^{-4}({\textstyle\frac{3}{8}}v^{4}+{\textstyle\frac{5}{2}}v^{2}w_{0}-4(\boldsymbol{\mathrm{v}}\cdot\boldsymbol{\mathrm{w}})+{\textstyle{\textstyle{\textstyle{1\over 2}}}}w_{0}^{2}-\Delta)\simeq 1.37\times 10^{-16}.$ As such, these terms are too small to account for GNSS applications in cislunar space, thus, they may be neglected.
Taking into account the relationship between TCB and TDB, given by (11), we develop an expression relating the rates of $\tt TDB$ and observer’s proper time, which from $d{\tt TDB}/{d\tau}=(d{\tt TDB}/d{\tt TCB})\,d{\tt TCB}/d\tau=\big(1-L_{\tt B}\big)\,d{\tt TCB}/d\tau$, results in:
$\frac{d{\tt TDB}}{d\tau}=(1-L_{\tt B})\Big(1+{1\over c^{2}}\Big\{{\textstyle\frac{1}{2}}v^{2}_{0}+\sum_{\tt B}{GM_{\tt B}\over r_{\tt B0}}\Big(1-J_{\tt B2}\Big(\frac{R_{\tt B}}{r_{\tt B0}}\Big)^{2}P_{20}(\cos\theta_{\tt B})\Big)\Big\}_{\tt\tau}\Big)+{\cal O}\Big(1.37\times 10^{-16}\Big).$ (58)
Integrating (58) over $\tau$ from $\tau_{0}$ to $\tau$ provides a relation between the proper time and $\tt TDB$:
${\tt TDB}=\big(1-L_{\tt B}\big)\Big(\Delta\tau+\frac{1}{c^{2}}\int^{\tau}_{\tau_{0}}\Big\{{\textstyle\frac{1}{2}}v^{2}_{0}+\sum_{\tt B}{GM_{\tt B}\over r_{\tt B0}}\Big(1-J_{\tt B2}\Big(\frac{R_{\tt B}}{r_{\tt B0}}\Big)^{2}P_{20}(\cos\theta_{\tt B})\Big)\Big\}_{\tau}dt\Big)+{\cal O}\Big(1.37\times 10^{-16}\Big)\Delta\tau,$ (59)
where the relativistic time equation at the clock is given as
$\tau=t^{*}-{1\over c^{2}}\int_{t_{0}}^{t^{*}}\Big\{{\textstyle\frac{1}{2}}v^{2}_{0}+\sum_{\tt B}{GM_{\tt B}\over r_{\tt B0}}\Big(1-J_{\tt B2}\Big(\frac{R_{\tt B}}{r_{\tt B0}}\Big)^{2}P_{20}(\cos\theta_{\tt B})\Big)\Big\}dt+{\cal O}\Big(1.10\times 10^{-16}\Big)(t^{*}-t_{0}),$ (60)
with the representation of the motion of the clock with respect to the BCRS given as below
$\bar{\boldsymbol{\mathrm{x}}}_{0}(\tau)=\boldsymbol{\mathrm{x}}_{0}(t^{*}),$ (61)
that may be solved either numerically or analytically, in the analogy to the approach used in (3)–(4). For direct transformations (56), one should integrate the integral either numerically or analytically. For the inverse transformation (59), to express the right-hand members in terms of $\tau$, one must obtain the function $t^{*}=t^{*}(\tau)$ by inverting the equation (60), also either numerically or analytically.
III Processing the GNSS data from the BCRS
In this Section, we discuss the validation of the position, velocity, and acceleration transformations outlined previously, now specifically tailored for GNSS spacecraft. Our focus is on evaluating the precision of spacecraft positioning achieved by processing data within the BCRS. This analysis aims to demonstrate the transformation’s effectiveness in enhancing the accuracy of GNSS spacecraft navigation.
Before proceeding, we will present the equations of motion for both coordinate systems and discuss the relevant dynamics. This overview is crucial for understanding how these transformations impact the positioning and navigation.
III.1 Equations of motion in the scaled TDB-compatible BCRS
III.1.1 BCRS metric in the mass monopole approximation
For many applications, it suffices to consider only the mass monopoles of all solar system bodies, retaining solely the masses which represents the mass monopole approximation of the BCRS metric, for a system of $N$ point-like gravitational sources in four dimensions is outlined in [1]:
$g_{00}$ $=$ $1-\frac{2}{c^{2}}\sum_{j\not=i}\frac{\mu_{j}}{r_{ij}}+\frac{2\beta}{c^{4}}\Big[\sum_{j\not=i}\frac{\mu_{j}}{r_{ij}}\Big]^{2}-\frac{1+2\gamma}{c^{4}}\sum_{j\not=i}\frac{\mu_{j}{\dot{r}}^{2}_{j}}{r_{ij}}+\frac{2(2\beta-1)}{c^{4}}\sum_{j\not=i}\frac{\mu_{j}}{r_{ij}}\sum_{k\not=j}\frac{\mu_{k}}{r_{jk}}-\frac{1}{c^{4}}\sum_{j\not=i}\mu_{j}\frac{\partial^{2}r_{ij}}{\partial t^{2}}+{\cal O}(c^{-5}),$
$g_{0\alpha}$ $=$ $\frac{2(\gamma+1)}{c^{3}}\sum_{j\not=i}\frac{\mu_{j}{\dot{\bf r}}^{\alpha}_{j}}{r_{ij}}+{\cal O}(c^{-5}),\qquad g_{\alpha\beta}=\gamma_{\alpha\beta}\Big(1+\frac{2\gamma}{c^{2}}\sum_{j\not=i}\frac{\mu_{j}}{r_{ij}}\Big)+{\cal O}(c^{-4}),$ (62)
where the indices $j$ and $k$ refer to the $N$ bodies, with $k$ including body $i$, whose motion is being investigated. The gravitational constant for body $j$, denoted as $\mu_{j}$, is defined as $\mu_{j}=Gm_{j}$, where $G$ is the universal Newtonian gravitational constant, and $m_{j}$ is the isolated rest mass of body $j$. Additionally, the vector $\boldsymbol{\mathrm{r}}_{i}$ represents the barycentric radius-vector of this body. The vector $\boldsymbol{\mathrm{r}}_{ij}=\boldsymbol{\mathrm{r}}_{\tt J}-\boldsymbol{\mathrm{r}}_{i}$ points from body $i$ to body $j$, $r_{ij}=|\boldsymbol{\mathrm{r}}_{\tt J}-\boldsymbol{\mathrm{r}}_{i}|$, and the vector $\boldsymbol{\mathrm{n}}_{ij}=\boldsymbol{\mathrm{r}}_{ij}/r_{ij}$ is the unit vector along this direction.
Furthermore, in (III.1.1), we have reinstated the parameterized post-Newtonian (PPN) parameters [31], which have been adapted to relativistic reference frames in [32]. The two PPN parameters, $\gamma$ and $\beta$, have specific physical interpretations: $\gamma$ quantifies the curvature of space-time produced by a unit rest mass, while $\beta$ quantifies the non-linearity in the superposition of gravitational fields in gravity theory. In general relativity, $\gamma=\beta=1$, and the other eight parameters are absent, situating the theory within a two-dimensional theoretical framework.
III.1.2 Equations of motion in BCRS: point-mass approximation
The barycentric metric given by (III.1.1) may be used to derive the post-Newtonian EIH equations of motion for a system of mass monopoles. These equations form the basis of modern solar system ephemerides. The relevant point-mass Newtonian and relativistic perturbative accelerations in the TDB-compatible BCRS are used to describe the motion of spacecraft in the solar system, including Earth orbiters and the Moon. The corresponding EIH equations are given as follows [1]:
$\ddot{\bf r}_{i}$ $=$ $\sum_{j\not=i}\frac{\mu_{j}({\bf r}_{j}-{\bf r}_{i})}{r_{ij}^{3}}\bigg\{1-\frac{2(\beta+\gamma)}{c^{2}}\sum_{l\not=i}\frac{\mu_{l}}{r_{il}}-\frac{2\beta-1}{c^{2}}\sum_{k\not=j}\frac{\mu_{k}}{r_{jk}}+\gamma\Big(\frac{{\dot{r}}_{i}}{c}\Big)^{2}+(1+\gamma)\Big(\frac{{\dot{r}}_{j}}{c}\Big)^{2}-\frac{2(1+\gamma)}{c^{2}}(\dot{\bf r}_{i}\cdot\dot{\bf r}_{j})$- (63)
$-~\frac{3}{2c^{2}}\Big[\frac{({\bf r}_{i}-{\bf r}_{j})\cdot{\dot{\bf r}}_{j}}{r_{ij}}\Big]^{2}+\frac{1}{2c^{2}}\big(({\bf r}_{j}-{\bf r}_{i})\cdot{\ddot{\bf r}}_{j}\big)\bigg\}+\frac{1}{c^{2}}\sum_{j\not=i}\frac{\mu_{j}}{r_{ij}^{3}}\Big\{\big({\bf r}_{i}-{\bf r}_{j}\big)\cdot\Big((2+2\gamma){\dot{\bf r}}_{i}-(1+2\gamma){\dot{\bf r}}_{j}\Big)\Big\}({\dot{\bf r}}_{i}-{\dot{\bf r}}_{j})+$
$+~\frac{3+4\gamma}{2c^{2}}\sum_{j\not=i}\frac{\mu_{j}{\ddot{\bf r}}_{j}}{r_{ij}}+\sum_{m=1}^{3}\frac{\mu_{m}({\bf r}_{m}-{\bf r}_{i})}{r^{3}_{im}}+\sum_{c,s,m}{\bf F}+{\cal O}(c^{-4}).$
The first term in (63) represents the Newtonian gravitational acceleration, which must be computed using barycentric TDB-compatible quantities. The subsequent terms are the post-Newtonian contributions. The penultimate term on the right-hand side of (63) refers to the ‘Big3’ asteroids—Ceres, Pallas, and Vesta—using the index mm. The final term accounts for forces on the Earth, Moon, and Mars from 297 other asteroids, which are grouped into three taxonomic classes (C, S, M), see [33] for details. For a planetary orbiter, it is usual to integrate the difference between the orbiter’s acceleration and the planet’s acceleration, both derived from the EIH equations. Typically, only the relativistic corrective terms with the index $j$ matching either the central planet or the Sun have significant effects.
III.1.3 Light time equation
To determine the orbits of planets and spacecraft, one must also account for the propagation of EM signals between any two points in space. The corresponding light-time equation is derived from the metric tensor in (III.1.1) as follows:
$t_{2}-t_{1}=\frac{r_{12}}{c}+(1+\gamma)\sum_{\tt B}\frac{GM_{\tt B}}{c^{3}}\ln\left[\frac{r_{1}^{\tt B}+r_{2}^{\tt B}+r_{12}^{\tt B}}{r_{1}^{\tt B}+r_{2}^{\tt B}-r_{12}^{\tt B}}\right]+{\cal O}(c^{-5}),$ (64)
where $t_{1}$ refers to the signal transmission time, and $t_{2}$ refers to the signal reception time. Additionally, $r_{1}$ and $r_{2}$ are the barycentric positions of the transmitter and receiver, respectively, while $r_{12}$ represents their spatial separation.
The Shapiro time delay is proportional to $(1+\gamma)$. The average solar potential at a distance of 1 AU is $\left<U\right>/c^{2}=0.99\times 10^{-8}$. The mean delay caused by the Sun for GPS signals is $\sim 2.66$ ns, equivalent to $\sim 79.8$ cm. The annual variation is $\pm 7$ mm. The delay due to Earth’s gravity can reach up to 42.3 ps (or $\sim 1.3$ cm), and it varies with the elevation angle. Therefore, only the Sun and Earth must be included in the model (64). At L-band, solar plasma delay is negligible except near conjunction; apply standard transformations for troposphere/ionosphere in the observable model. Most of the solar term behaves as an overall scale in range and cancels in differenced observables, but should be retained for absolute light-time consistency at the centimeter level.
For increased fidelity near conjunctions, evaluate the Shapiro term in (64) with retarded ephemeris positions of gravitating bodies $\tt B$:
$\mathbf{x}_{\tt B}^{\mathrm{ret}}(t_{i})\simeq\mathbf{x}_{\tt B}(t_{i})-\frac{r_{\tt B,\rm i}}{c}\,\mathbf{v}_{\tt B}(t_{i}),\quad r_{\tt B,\rm i}\equiv\bigl\|\mathbf{r}_{i}-\mathbf{x}_{\tt B}^{\mathrm{ret}}(t_{i})\bigr\|,\quad i\in\{1,2\}.$
This accounts for leading $\mathcal{O}(v_{\tt B}/c)$ corrections without changing the observable interface. The light-time and its derivatives used here are consistent with the time-transfer-function (TTF) formalism; see e.g., [34] for the moving-axisymmetric-body generalization used in interplanetary radioscience.
When either endpoint lies on the rotating Earth, the one-way coordinate light-time acquires the Sagnac correction due to the rotation vector $\boldsymbol{\Omega}_{\oplus}$,
$\Delta t_{\rm sag}\;=\;\frac{1}{c^{2}}\big(\boldsymbol{\Omega}_{\oplus}\cdot\big[\mathbf{r}_{\rm R}(t_{\rm r})\times\mathbf{r}_{\rm T}(t_{\rm e})\big]\big),$ (65)
so the total is $\Delta t=\Delta t_{\rm geom}+\Delta t_{\rm Shapiro}+\Delta t_{\rm sag}$, with $\Delta t_{\rm geom}$ and $\Delta t_{\rm Shapiro}$ as in (64). Numerically, for a typical GNSS MEO geometry with $|\mathbf{r}_{\rm R}|\!\approx\!6378~\mathrm{km}$, $|\mathbf{r}_{\rm T}|\!\approx\!26{,}560~\mathrm{km}$, and $|\boldsymbol{\Omega}_{\oplus}|\!=\!7.292115\times 10^{-5}\,\mathrm{s^{-1}}$, one finds $|\Delta t_{\rm sag}|\!\lesssim\!1.4\times 10^{-7}\,\mathrm{s}$ ($\approx\!41~\mathrm{m}$ in range), hence mandatory for centimeter modeling of ground links [35]. The term (65) leads to a correction
$\frac{\partial}{\partial t_{\rm r}}\Delta t_{\rm sag}=\frac{1}{c^{2}}\Big(\boldsymbol{\Omega}_{\oplus}\cdot\Big(\big[\,\mathbf{v}_{\rm R}(t_{\rm r})\times\mathbf{r}_{\rm T}(t_{\rm e})\big]-\big[\mathbf{v}_{\rm T}(t_{\rm e})\times\mathbf{r}_{\rm R}(t_{\rm r})\big]\Big)\Big),$ (66)
that is required wherever one forms two‑way frequency/phase models. For a rotating ground site this term is at the $10^{-12}-10^{-11}$ fractional level for L-band.
Frequency transfer observable at $\mathcal{O}(c^{-3})$
For one-way transfer from emitter A to receiver B,
$\frac{\nu_{\tt A}}{\nu_{\tt B}}=\frac{(u^{0})_{\tt A}}{(u^{0})_{\tt B}}\,\frac{q_{\tt A}}{q_{\tt B}}+\mathcal{O}(c^{-4}),$ (67)
where the ratio $q_{A}/q_{B}$ follows from differentiating the light-time (our Eq. (64)) with respect to emission time, retaining terms through $\mathcal{O}(c^{-3})$ [36, 37]. For the two-way (uplink+downlink) configuration used in precise time/frequency comparisons, a convenient form is
$\frac{\nu_{\tt B}}{\nu_{\tt A}}=\frac{1}{2}\left(\frac{\nu_{\tt B}}{\nu_{\tt B^{\prime}}}\right)_{\!\text{station}}+\Delta_{\tt AB}+\frac{1}{2}+\mathcal{O}(c^{-4}),$ (68)
with
$\Delta_{\tt AB}$ $=\frac{1}{c^{2}}\!\left(U_{\tt AB}-{\textstyle\frac{1}{2}}v_{\tt AB}^{2}-(\mathbf{R}_{\tt AB}\cdot\mathbf{a}_{\tt B})\right)\!\left(1+\frac{(\mathbf{N}_{\tt AB}\cdot\mathbf{v}_{\tt AB})}{c}\right)-\frac{R_{\tt AB}}{c^{3}}\Big((\mathbf{v}_{\tt A}\cdot\mathbf{a})_{\tt B}-R_{\tt AB}\,b_{\tt B}-2(\mathbf{v}_{\tt B}\cdot\mathbf{a}_{\tt B})+(\mathbf{v}_{\tt B}\cdot\nabla U_{\tt B})\Big),$ (69)
where $U_{\tt AB}=U_{\tt B}-U_{\tt A}$, $\mathbf{R}_{\tt AB}=\mathbf{r}_{\tt B}-\mathbf{r}_{\tt A}$, $R_{\tt AB}=|\mathbf{R}_{\tt AB}|$, $\mathbf{N}_{\tt AB}=\mathbf{R}_{\tt AB}/R_{\tt AB}$, $v_{\tt AB}=\mathbf{v}_{\tt A}-\mathbf{v}_{\tt B}$, and $b_{\tt B}=d(\mathbf{N}_{\tt AB}\cdot\mathbf{a}_{\tt B})/dt$. These expressions are consistent with our light-time Eq. (64) and close the model for range-rate and frequency transfer at the level needed for $10^{-16}$ work.
III.2 Equations of motion in the scaled TT-compatible GCRS
III.2.1 GCRS: Practically-relevant formulation
In practical GCRS implementations, one includes all post-Newtonian terms up to $\mathcal{O}(c^{-4})$ in $G_{00}$, $\mathcal{O}(c^{-3})$ in $G_{0\alpha}$, and $\mathcal{O}(c^{-2})$ in $G_{\alpha\beta}$, but discards any metric perturbations smaller than $5\times 10^{-18}$. Accordingly, the metric tensor of GCRS (see [9, 19]) retaining only $\lvert\delta G_{mn}\rvert\gtrsim 5\times 10^{-18}$, sufficient for high-precision time-keeping applications, becomes
$G_{00}(T,{\bf X})$ $=$ $1-\frac{2}{c^{2}}\Big\{W_{\tt E}(T,{\bf X})+W_{\rm tid}(T,{\bf X})\Big\}+\frac{2}{c^{4}}W^{2}_{\rm E}(T,{\bf X})+{\cal O}\Big(c^{-5};6.61\times 10^{-25}\Big),$ (70)
$G_{0\alpha}(T,{\bf X})$ $=$ $-{2G\over c^{3}}\frac{[{\boldsymbol{\mathrm{S}}}_{\tt E}\times{\boldsymbol{\mathrm{X}}}]_{\alpha}}{R^{3}}+{\cal O}\Big(c^{-5};2.79\times 10^{-19}\Big),$ (71)
$\ G_{\alpha\beta}(T,{\bf X})$ $=$ $\gamma_{\alpha\beta}\Big(1+\frac{2}{c^{2}}\Big\{W_{\tt E}(T,{\bf X})+W_{\rm tid}(T,{\bf X})\Big\}\Big)+{\cal O}\Big(c^{-4};5.57\times 10^{-20}\Big),$ (72)
where $S_{\tt E}\simeq 5.86\times 10^{33}\,{\rm kg\,m}^{2}$/s is the Earth spin vector moment. Also, $W_{\tt E}(T,{\bf X})$ is the Earth’s gravitational potential
$W_{\tt E}(T,{\bf X})$ $=$ ${GM_{\tt E}\over R}\Big\{1+\sum_{\ell=2}^{\infty}\sum_{m=0}^{\ell}\Big({R_{\tt E}\over R}\Big)^{\ell}P_{\ell m}(\cos\theta)\Big(C^{\tt E}_{\ell m}(T,R)\cos m\phi+S^{\tt E}_{\ell m}(T,R)\sin m\phi\Big)\Big\}+{\cal O}(c^{-4}),$ (73)
where $M_{\tt E}$ and $R_{\tt E}$ are the Earth’s mass and equatorial radius, respectively, while $P_{\ell k}$ are the associated Legendre-polynomials [38]. $W_{\rm tid}(T,{\bf X})$ is the Newtonian the tidal potential by external bodies
$W_{\rm tid}(T,{\bf X})=U_{\rm ext}({\bf x}_{\tt E}+{\bf X})-U_{\rm ext}({\bf x}_{\tt E})-\big({\bf X}\cdot\boldsymbol{\mathrm{\nabla}}U_{\rm ext}({\bf x}_{\tt E})\big),$ (74)
[…]
The error bounds in (70)–(72) are due to the dominant omitted corrections evaluated at GPS altitude, specifically: $\delta G_{00}^{\rm(mix)}=-4c^{-4}W_{\tt E}\,W_{\rm tid}\simeq 4c^{-4}(GM_{\tt E}/r_{\tt GPS})(GM_{\tt S}/{\rm AU}^{3}+Gm_{\tt M}/r_{\tt EM}^{3})r_{\tt GPS}^{2}\simeq 6.61\times 10^{-25}$; $\delta G_{0\alpha}^{\rm(tid)}=-4c^{-3}W_{\rm tid}^{\alpha}\simeq 4c^{-3}(GM_{\tt M}/r_{\tt EM}^{3})v_{\tt M}r_{\tt GPS}^{2}\simeq 2.79\times 10^{-19}$; and $\delta G_{\alpha\beta}^{\rm(2PN)}=\gamma_{\alpha\beta}\,\tfrac{3}{2}c^{-4}W_{\tt E}^{2}\simeq\tfrac{3}{2}c^{-4}(GM_{\tt E}/r_{\tt GPS})^{2}\simeq 4.78\times 10^{-20}$, see details in [19]. Although we evaluated the metric components in (70)–(72) at GPS altitude (i.e., where tidal contributions exceed those at the surface), these expressions remain valid for all Earth‐orbit regimes from MEO through GEO.
III.2.2 Equations of motion in the GCRS
The trajectory of an Earth satellite in GCRS is governed by the following equation of motion [30, 1, 39, 15]:
$\ddot{\boldsymbol{\mathrm{r}}}(t)$ $=$ $\ddot{\boldsymbol{\mathrm{r}}}_{\tt E}+\ddot{\boldsymbol{\mathrm{r}}}_{{\tt E}J_{2}}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt tidal}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt GR}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt SRP}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt{\tt E}elast}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt{\tt E}rad}+\ddot{\boldsymbol{\mathrm{r}}}_{\tt custom},$ (75)
where $\ddot{\boldsymbol{\mathrm{r}}}_{\tt E}$ is the Newtonian acceleration due to the gravity field of the Earth [^2] Note that, for convenience, we are using lowercase letters to denote vectors in the GCRS, rather than uppercase letters.; $\ddot{\boldsymbol{\mathrm{r}}}_{{\tt E},J_{2}}$ is the acceleration due to Earth’s oblateness; $\ddot{\boldsymbol{\mathrm{r}}}_{\tt tidal}$ is the acceleration due to the tidal gravity field induced by the Sun, Moon and Venus; $\ddot{\boldsymbol{\mathrm{r}}}_{\tt GR}$ is the acceleration due to general relativity; $\ddot{\boldsymbol{\mathrm{r}}}_{\tt SRP}$ is the acceleration due to direct/indirect solar radiation pressure; $\ddot{\boldsymbol{\mathrm{r}}}_{\rm{\tt E}elast}$ is the acceleration due to elastic Earth response, pole, and ocean tides; $\ddot{\boldsymbol{\mathrm{r}}}_{\rm{\tt E}rad}$ is the acceleration due to Earth albedo and infra-red (IR) radiation; and $\ddot{\boldsymbol{\mathrm{r}}}_{\rm custom}$ is the acceleration due to any custom force model. Note that all quantities (positions, accelerations, distances, mass parameters) in (75) are TT-compatible.
The acceleration of a spacecraft at $\boldsymbol{\mathrm{r}}$, due to the Earth’s gravity field in (75) is computed as usual
$\ddot{\boldsymbol{\mathrm{r}}}_{\tt E}={\boldsymbol{\mathrm{\nabla}}}U_{\tt E}(\boldsymbol{\mathrm{r}})+{\cal O}(c^{-4})$ $=$ $-\frac{GM_{\tt E}}{r^{3}}{\boldsymbol{\mathrm{r}}}\Big(1+{\cal O}\Big(\frac{R^{2}_{\tt E}}{r^{2}}\Big)\Big)\simeq-\frac{GM_{\tt E}}{a^{2}}{\boldsymbol{\mathrm{n}}}_{r}\approx 0.57~{\rm m/s}^{2},$ (76)
where $\boldsymbol{\mathrm{a}}=a\boldsymbol{\mathrm{n}}_{r}$ is the semi-major axis of the spacecraft orbit, ${\boldsymbol{\mathrm{n}}}_{r}=\boldsymbol{\mathrm{r}}/r$ is the unit vector in the direction of the spacecraft from the GCRS. In this estimate, we truncated the Earth relativistic gravitational potential $U_{\tt E}(\boldsymbol{\mathrm{x}})\equiv W_{\tt E}(T,{\bf X})$ from (73) to the monopole term. Clearly, this is the dominant term in the equations of motion of an Earth-orbiting satellite (75).
Let $\mu_{\tt E}=GM_{\tt E}$, $R_{\tt E}$ the Earth’s mean equatorial radius, $J_{2,\tt E}$ the quadrupole coefficient, $\hat{\mathbf{r}}=\mathbf{r}/r$, and $\hat{\mathbf{S}}$ the unit spin axis of the Earth. The 1PN gravitoelectric acceleration due to the oblateness is
$\ddot{\boldsymbol{\mathrm{r}}}_{{\tt E},J_{2}}=$ $\frac{3J_{2,\tt E}\,\mu_{\tt E}R_{\tt E}^{2}}{2\,c^{2}\,r^{4}}\Big(\big[5\,\hat{\mathbf{r}}(\hat{\mathbf{S}}\!\cdot\!\hat{\mathbf{r}})^{2}-2\,\hat{\mathbf{S}}(\hat{\mathbf{S}}\!\cdot\!\hat{\mathbf{r}})-\hat{\mathbf{r}}\big]\big[v^{2}-\frac{4\mu_{\oplus}}{r}\big]$-
[…]
In the worst-case orientation and using $J_{2,\oplus}=1.08263\times 10^{-3}$, $R_{\oplus}=6378~\mathrm{km}$, $\mu_{\tt E}=3.986\times 10^{14}~\mathrm{m^{3}s^{-2}}$, this acceleration was evaluated to contribute: LEO: $\sim 7\times 10^{-11}$ m s-2, MEO: $\sim 10\times 10^{-13}$ m s-2, cislunar: $\sim 5\times 10^{-16}$ m s-2. Thus, for LEO its omission biases precise orbit and gravity solutions; for MEO it is small but systematic; for cislunar it is negligible for dynamics but enter clock/frequency transfers only via $U$ coupling already modeled.
The tidal acceleration due to the gravity fields of the Sun, Venus and Moon are also well known. Using $u^{\tt tidal}_{\tt E}(\boldsymbol{\mathrm{x}})\equiv W_{\tt tidal}(\boldsymbol{\mathrm{x}})$ from (74), we compute it as usual:
$\ddot{\boldsymbol{\mathrm{r}}}_{\rm tidal}$ $=$ ${\boldsymbol{\mathrm{\nabla}}}u^{\tt tidal}_{\tt E}(\boldsymbol{\mathrm{x}})+{\cal O}\Big(\frac{x^{3}}{r^{4}_{\tt BE}},c^{-2}\Big)\simeq\sum_{b\not={\rm E}}\frac{GM_{b}}{r^{3}_{\tt BE}}\Big(3(\boldsymbol{\mathrm{n}}_{\tt BE}\cdot\boldsymbol{\mathrm{r}})\boldsymbol{\mathrm{n}}_{\tt BE}-\boldsymbol{\mathrm{r}}\Big)\leq$ (78)
$\leq$ $\frac{2GM_{\tt S}a}{a^{3}_{\tt E}}(\boldsymbol{\mathrm{n}}_{\tt S}\cdot\boldsymbol{\mathrm{n}}_{r})\boldsymbol{\mathrm{n}}_{\tt S}+\frac{2GM_{\tt M}a}{a^{3}_{\tt M}}(\boldsymbol{\mathrm{n}}_{\tt M}\cdot\boldsymbol{\mathrm{n}}_{r})\boldsymbol{\mathrm{n}}_{\tt M}+\frac{2GM_{\tt V}a}{a^{3}_{\tt V}}(\boldsymbol{\mathrm{n}}_{\tt V}\cdot\boldsymbol{\mathrm{n}}_{r})\boldsymbol{\mathrm{n}}_{\tt V}\approx$
$\approx$ $2.11\times 10^{-6}~{\rm m/s}^{2}+4.59\times 10^{-6}~{\rm m/s}^{2}+2.45\times 10^{-10}~{\rm m/s}^{2},$
where $a_{\tt E},e_{\tt E}$, and ${\cal E}_{\tt E}$ and $a_{\tt V},e_{\tt V}$, and ${\cal E}_{\tt V}$ are the semi-major axis, eccentricity and eccentric anomaly characterizing the heliocentric orbits of the Earth and Venus correspondingly, while $a_{\tt M},e_{\tt M}$, and ${\cal E}_{\tt M}$, are the same quantities for the geocentric orbit of the Moon. We will use the value of $a_{\tt M}=384,400$ km to denote the Earth-Moon mean distance.
In addition, we estimate the maximal accelerations from the tides introduced by Jupiter and Saturn:
$|\ddot{\boldsymbol{\mathrm{r}}}^{\tt J}_{\rm tidal}|$ $\lesssim$ $\frac{2GM_{\tt J}a}{(a_{\tt J}-a_{\tt E})^{3}}\approx 2.70\times 10^{-11}~{\rm m/s}^{2},\qquad|\ddot{\boldsymbol{\mathrm{r}}}^{\tt Sa}_{\rm tidal}|\lesssim\frac{2GM_{\tt Sa}a}{(a_{\tt Sa}-a_{\tt E})^{3}}\approx 9.52\times 10^{-13}~{\rm m/s}^{2},$ (79)
with contributions from Uranus, Neptune being below $10^{-14}$ m/s2. Note that these tidal forces are not accounted for in the equations of motion used to process GPS observables. Due to the long orbital periods of these planets, they contribute a nearly constant background that is removed along with other constant terms during data analysis.
The relativistic corrections to the accelerations of Earth satellites written in a local inertial GCRS that is kinematically non-rotating read as [40]:
$\ddot{\boldsymbol{\mathrm{r}}}_{\tt GR}$ $=$ $\frac{GM_{\tt E}}{c^{2}r^{3}}\bigg\{\Big[2(\beta+\gamma)\frac{GM_{\tt E}}{r}-\gamma\dot{\boldsymbol{\mathrm{r}}}^{2}\Big]\boldsymbol{\mathrm{r}}+2(1+\gamma)({\boldsymbol{\mathrm{r}}}\cdot\dot{\boldsymbol{\mathrm{r}}})\dot{\boldsymbol{\mathrm{r}}}\bigg\}+$ (80)
$+\,(1+\gamma)\frac{GM_{\tt E}}{c^{2}r^{3}}\Big[\frac{3}{r^{2}}[\boldsymbol{\mathrm{r}}\times\dot{\boldsymbol{\mathrm{r}}}](\boldsymbol{\mathrm{r}}\cdot\boldsymbol{\mathrm{S}}_{\tt E})+[\dot{\boldsymbol{\mathrm{r}}}\times\boldsymbol{\mathrm{S}}_{\tt E}]\Big]+(1+2\gamma)\frac{GM_{\tt S}}{c^{2}r^{3}_{\tt E}}\big[[{\boldsymbol{\mathrm{r}}}_{\tt E}\times\dot{\boldsymbol{\mathrm{r}}}_{\tt E}]\times\dot{\boldsymbol{\mathrm{r}}}\big],$
where $GM_{\tt E}$ is the standard Earth gravitation product; $GM_{\tt S}$ is the standard Sun gravitation product; $\beta,\gamma$ are PPN parameters equal to 1 in GR; $c$ is the speed of light; $\boldsymbol{\mathrm{r}},\dot{\boldsymbol{\mathrm{r}}}$ are the position and velocity of a satellite with respect to the geocenter; $\boldsymbol{\mathrm{r}}_{\tt E},\dot{\boldsymbol{\mathrm{r}}}_{\tt E}$ are the position and velocity of the Earth with respect to the Sun; ${\boldsymbol{\mathrm{S}}}_{\tt E}$ is the Earth’s spin moment (angular momentum per unit of mass). Introducing $\mathcal{R}$ as the rotation matrix transformations ECEF $\to$ inertial, we write $S_{\tt E}\simeq 0.33068\,R_{\tt E}^{2}\,\omega_{\tt E}\,\mathcal{R}\,\hat{\mathbf{z}}$.
The first line describes the so-called Schwarzschild term [41, 42], the next term is the frame-dragging gravitomagnetic Lense-Thirring effect [43], whereas the third corresponds to the de Sitter effect [44]. Other relativistic effects are related to, e.g., other celestial bodies, which can be approximated as small corrections to the non-relativistic tidal forces, Thomas precession, or Earth’s oblateness quadruple term; however, they are currently not considered in the precise GNSS orbit solutions, due to their very small magnitudes [30, 40, 18, 45, 39].
We evaluate the Schwarzschild term in (80) as:
$\ddot{\boldsymbol{\mathrm{r}}}_{\rm Sch}$ $=$ $\frac{GM_{\tt E}}{c^{2}r^{3}}\bigg\{\Big[2(\beta+\gamma)\frac{GM_{\tt E}}{r}-\gamma\dot{\boldsymbol{\mathrm{r}}}^{2}\Big]\boldsymbol{\mathrm{r}}+2(1+\gamma)({\boldsymbol{\mathrm{r}}}\cdot\dot{\boldsymbol{\mathrm{r}}})\dot{\boldsymbol{\mathrm{r}}}\bigg\}\simeq\frac{3(GM_{\tt E})^{2}}{c^{2}a^{3}}{\textstyle\frac{1}{3}}(2\beta+\gamma)\boldsymbol{\mathrm{n}}_{r}\leq 2.83\times 10^{-10}~{\rm m/s}^{2}.$~~~ (81)
Next, we evaluate the frame-dragging gravitomagnetic Lense-Thirring effect in (80):
$\ddot{\boldsymbol{\mathrm{r}}}_{\tt LT}$ $=$ $(1+\gamma)\frac{GM_{\tt E}}{c^{2}r^{3}}\Big[\frac{3}{r^{2}}[\boldsymbol{\mathrm{r}}\times\dot{\boldsymbol{\mathrm{r}}}](\boldsymbol{\mathrm{r}}\cdot\boldsymbol{\mathrm{S}}_{\tt E})+[\dot{\boldsymbol{\mathrm{r}}}\times\boldsymbol{\mathrm{S}}_{\tt E}]\Big]\simeq 1.98\Big(\frac{GM_{\tt E}}{a}\Big)^{\frac{3}{2}}\frac{R^{2}_{\tt E}\omega_{\tt E}}{c^{2}a^{2}}\,{\textstyle\frac{1}{2}}\sin 2i_{0}\sin i_{\tt ecl}\leq 0.98\times 10^{-12}~{\rm m/s}^{2},$~~~~~~ (82)
where we used the unit vectors ${\boldsymbol{\mathrm{n}}}_{[\boldsymbol{\mathrm{r}}\times\dot{\boldsymbol{\mathrm{r}}}]}=[\boldsymbol{\mathrm{r}}\times\dot{\boldsymbol{\mathrm{r}}}]/|\boldsymbol{\mathrm{r}}\times\dot{\boldsymbol{\mathrm{r}}}|$, $\boldsymbol{\mathrm{n}}_{r}=\boldsymbol{\mathrm{r}}/r$ and $\boldsymbol{\mathrm{n}}_{v}=\boldsymbol{\mathrm{v}}/v$. In addition, we introduce the unit vector ${\boldsymbol{\mathrm{n}}}^{\prime}_{z}=\hat{\boldsymbol{\mathrm{T}}}{\boldsymbol{\mathrm{n}}}_{z}$ relating the Earth’s angular momentum with the direction to the north pole. Also, $i_{0}$ is the spacecraft’s orbital inclination, $i_{0}=55^{\circ}$ for GPS orbits, and where $i_{\tt ecl}=23.5^{\circ}$ is Earth’s axial tilt.
Finally, we consider the de Sitter effect due to the Sun in (80):
$\ddot{\boldsymbol{\mathrm{r}}}^{\tt S}_{\tt dS}$ $=$ $(1+2\gamma)\frac{GM_{\tt S}}{c^{2}r^{3}_{\tt E}}\big[[{\boldsymbol{\mathrm{r}}}_{\tt E}\times\dot{\boldsymbol{\mathrm{r}}}_{\tt E}]\times\dot{\boldsymbol{\mathrm{r}}}\big]\simeq\frac{3}{c^{2}a_{\tt E}}\Big(\frac{GM_{\tt S}}{a_{\tt E}}\Big)^{\frac{3}{2}}\sqrt{\frac{GM_{\tt E}}{a}}\,{\textstyle\frac{1}{3}}(1+2\gamma)\sin i_{\tt ecl}\leq 8.92\times 10^{-12}~{\rm m/s}^{2}.$ (83)
Thus, relativistic contributions are important and must be kept in the model.
III.2.3 Transforming equations of motion from GCRS to BCRS
The metric tensor (70)–(72) along with the gravitational potentials (73)–(74) describes spacetime in the GCRS, which we adopt to formulate the relativistic model for timing and frequency observables. Detailed derivations of the GCRS formulation can be found in [18, 12, 11, 13]. As a result, within the local GCRS, the following gravitational accelerations due to mass monopoles must be considered:
  • The Newtonian gravitational acceleration from the Earth, given by (76).
  • The relativistic correction to the Earth’s gravitational acceleration, provided by the leading term in (80).
  • The accelerations associated with De Sitter and Lense-Thirring precessions, given by the 2-nd and 3-rd terms in (80).
  • The third-body Newtonian acceleration due to tidal effects from all other solar system bodies, described by (78).
Accurately calculating third-body acceleration is challenging due to the limitations of the GCRS in modeling interactions with celestial bodies. The procedure, as outlined in [1], Sec. 4.5.1, involves the following steps:
  • Convert the integration time from TT to TDB along the spacecraft’s worldline. As the difference between TT and TDB is periodic and typically less than 2 ms, this correction is generally negligible.
  • Obtain the TDB-compatible BCRS state vectors for both the Earth and the third body from the planetary ephemeris at the corresponding TDB epoch.
  • Determine the position vectors from the spacecraft to the third body and from Earth to the third body using the TDB-compatible BCRS states of the Earth and the third body, combined with the TT-compatible GCRS state of the spacecraft.
  • Compute the third-body acceleration using these position vectors and the mass parameter of the third body from the BCRS, ensuring compatibility with the TDB time scale.
  • Apply a scaling factor to the computed third-body acceleration by multiplying it by $(1-L_{\tt C})$, where $L_{\tt C}$ accounts for the relativistic corrections. Implementation guard: Apply the $(1-L_{\tt C})$ factor only if the third-body term was built from TT-compatible constants. If you compute the term with TDB-compatible states and GM values (e.g., DE440), do not apply an additional $(1-L_{\tt C})$; the scaling in Sec. II.4 already accounts for it.
Note that relativistic corrections to third-body perturbations are typically small and often disregarded. The main effect is due to the Sun’s gravitational influence, with smaller contributions from other planets, causing the local inertial frames near Earth to experience De Sitter precession relative to the BCRS in the TDB frame. Nevertheless, the TT-compatible GCRS remains kinematically non-rotating relative to the TDB-compatible BCRS. This non-rotation creates a non-inertial reference frame, requiring the inclusion of Coriolis acceleration.
The De Sitter acceleration (83), is a relativistic correction to 3rd-body acceleration and is implicitly included when using the EIH equations in the BCRS (see [1], Sec. 4.4.2). It is important to note that Lense-Thirring precession (82) (see also [1], Secs. 4.4.3 and 4.5.4), is an independent effect caused by the angular momentum of the Earth (or a planet) and is not included in the EIH equations. The full set of relativistic correction terms is provided in (80).
III.2.4 Considering the gravitational harmonics of the Earth
The Earth’s motion around the Sun induces relativistic distortions in the gravitational spherical harmonics when observed in the BCRS. These distortions complicate the accurate computation of accelerations arising from these harmonics within the barycentric system. To mitigate these effects, it is necessary to apply relativistic corrections, which involve several key steps detailed in [1], Sec. 4.4.5:
  • Transform the spacecraft’s position vector from the TDB-compatible BCRS to the TT-compatible GCRS using a Lorentz transformation (refer to [1], Eq. (4-11)).
  • Calculate the acceleration due to gravitational spherical harmonics in the TT-compatible GCRS.
  • Convert the computed acceleration back to the barycentric frame using the method outlined in [1], Eq. (4-55).
This method can be applied to other planets as well. Given the relativistic distortions present in the barycentric reference frame, performing spherical harmonic expansions in a planetocentric system provides improved precision. Additionally, it is imperative to confirm that the reference radius and mass parameters used in the spherical harmonic expansion are properly scaled to account for relativistic effects. If these parameters do not reflect the appropriate relativistic scaling, they must be adjusted to ensure accuracy in the calculations.
This method accounts for the relativistic adjustments that arise from transformations between different space-time coordinate systems. Incorporating these corrections significantly improves the precision of the calculated acceleration due to the geopotential in the geocentric reference frame. Notably, the relativistic correction associated with the acceleration from Earth’s quadrupole moment, $J_{2}$, is extensively discussed in [46], Sec. 4.2.3.
III.3 Practical implementation and validation
In this Section, we outline the validation tests performed to verify the proper implementation of the EIH equations within the interplanetary orbit determination system used by GNSS spacecraft. The testing involved comparing the accelerations computed in the BCRS with those calculated in the GCRS. To enable this comparison, the accelerations derived in the GCRS were transformed into the BCRS using a series of relativistic transformations described in Sec. II.4. This approach ensures that the system accurately incorporates the relativistic effects necessary for precise navigation and tracking of GNSS spacecraft.
Figure 1: Millimeter‑level closure test of the BCRS implementation with JPL GipsyX. Panels show radial/along‑track/cross‑track closure residuals in meters (note the $10^{-3}$ scale, i.e., millimeter range) versus integration time (hours).
Table 4: Configuration used for the closure test shown in Fig. 1.
Item Value
Ephemeris JPL DE440 (argument: TDB)
Earth gravity Spherical harmonics to degree/order $N=M=20$ ($T$-compatible)
Third bodies Sun and Moon tidal terms (Newtonian)
Relativity (GCRS) Schwarzschild, de Sitter, Lense–Thirring per Eq. (80)
Relativity (BCRS) EIH point-mass terms per Eq. (63)
SRP/tides Standard GipsyX models (identical in both frames)
Span 24 h; start epoch UTC (specified in run log)
transformations Full vs. screened: Eqs. (36) [position], (42) [velocity], (51) [acceleration]
To validate the new relativistic framework we have developed a rigorous development and testing plan with GipsyX [17]. GipsyX is precision software developed by JPL for GNSS-based positioning, navigation, timing, and Earth science applications. It supports GPS, GLONASS, BeiDou, and Galileo systems and offers precise, centimeter-level positioning for various platforms, including low-Earth orbiters and terrestrial stations. GipsyX integrates real-time and post-processing capabilities and is continuously updated to meet the evolving demands of space geodesy. [^3] For further details on GipsyX, please visit https://gipsyx.jpl.nasa.gov/
We have undertaken a comprehensive implementation approach to enhance the accuracy of GNSS positioning by incorporating advanced computational models:
  • Imported and used the JPL planetary ephemeris DE440 for high-precision celestial mechanics.
  • Implemented the BCRS, integrating state-of-the-art relativistic models to ensure precise trajectory computations.
  • Developed multiple centers of integration to facilitate robust data handling and computational efficiency.
  • Integrated models for lunar orientation, lunar gravity, and lunar albedo to enhance lunar navigation capabilities.
To validate the implementation of the barycentric coordinate reference frame and the associated transformations, we conducted a series of rigorous tests:
  1. 1.
    Integrated a GPS orbit within the GCRS to establish a baseline for comparison.
  2. 2.
    Transformed the orbital state from the GCRS to the BCRS, while maintaining high accuracy in the transformation process.
  3. 3.
    Re-integrated the GPS orbit in the BCRS, starting from the initial condition transformed from the GCRS.
  4. 4.
    Compared the integrated orbit in the BCRS (from Step 3) with the transformed orbit in the GCRS (from Step 2) to assess the precision and accuracy of the transformations.
This testing procedure ensures that the implemented models and transformations meet the stringent requirements necessary for sub-millimeter accuracy in GNSS spacecraft navigation.
The numerical analysis conducted between the GCRS and the BCRS demonstrates that orbit states and observables can be precisely aligned, provided that the appropriate space-time transformations and relativistic corrections are applied. In the barycentric system, it is crucial not to overlook the relativistic corrections to acceleration. While in the Earth-centric system, these relativistic corrections to acceleration are less critical, the necessity for space-time transformations for conversion to and from the BCRS remains significant. In our testing scenario, after implementing the relativistic corrections and space-time transformations between the two systems, the discrepancy was observed to be under a millimeter in all three directions, see Fig. 1 within one day.
IV Conclusions and next steps
We conducted a comprehensive analysis of general relativistic effects on the dynamics of near-Earth satellites within the GCRS and BCRS frameworks. Our results demonstrate that, when full relativistic corrections are applied, both reference systems yield consistent and equivalent results, enabling high-precision satellite navigation. These corrections are essential for achieving sub-centimeter accuracy in orbit determination and picosecond-level precision in time synchronization, which are especially critical for long-duration and deep-space missions, e.g., [47]. Our analysis supports the use of GNSS signals for precise navigation and timing within the Earth-Moon system, laying the foundation for future cislunar operations where cm-level accuracy in spatial positioning is required over large spatial regions.
Upcoming cislunar missions will necessitate a similar relativistic approach to ensure accurate time and position transformations between the BCRS, Earth, and the Moon. This calls for the development of a Lunicentric Coordinate Reference System (LCRS), analogous to the GCRS (see discussion in Appendix A and [9, 19].) A key step will be establishing a TDB-compatible time scale for the Moon, such as TCL (Temps Coordonné Lunaire), to synchronize lunar timekeeping with the BCRS. Additionally, the spatial scale in lunar coordinates must incorporate relativistic effects, such as Lorentz contraction, to meet the sub-millimeter accuracy demands of future lunar missions.
[…]
The methodology used to develop and validate the relativistic GNSS navigation framework in GCRS and BCRS can be extended to the LCRS to achieve millimeter-level orbit closure and picosecond-level timing consistency in GipsyX. Looking ahead, we identify three potential pathways to mature the LCRS into a flight-ready capability:
  1. 1.
    Lunar GNSS Pathfinder Missions: Integrate a compact GNSS receiver and atomic clock payload (e.g. NASA’s Deep Space Atomic Clock DSAC–II or JPL’s Next-Gen On-Orbit Navigator) aboard a Commercial Lunar Payload Services (CLPS) lander or Lunar Gateway module to measure real-time LCRS time offsets and validate centimeter-scale positioning in situ.
  2. 2.
    Lunar Relay Satellite Demonstrations: Deploy a smallsat in a halo orbit about the Earth–Moon L2 point, equipped with dual-frequency GNSS antennas synchronized to a space‐qualified hydrogen maser. This will exercise the LCRS transformations under large Sun–Moon gravitational gradients and characterize residual frame-bias errors below 1 cm.
  3. 3.
    Clock‐Experiment Campaigns: Perform precision time‐transfer tests between DSAC-class clocks on Earth‐orbiting GNSS platforms and LCRS-referenced clocks on lunar orbiters, quantifying relativistic scale factors and assessing long-term stability at the $10^{-17}$ level using extended relativistic framework (e.g., [19].)
By executing these demonstrations—first in MEO, then on the lunar surface, and finally via spacecraft in cislunar space—we will close the loop on both navigation and timing, fully leveraging GNSS signals across the Earth–Moon system. This LCRS deployment roadmap paves the way for cm-accurate cislunar navigation and ps-synchronized operations, enabling reliable guidance for Artemis, commercial landers, and future deep-space exploration.
Ongoing work focuses on refining the relativistic corrections and coordinate transformations across GCRS, BCRS, and LCRS, and on integrating them into operational systems such as JPL’s GipsyX. These efforts are essential for upcoming missions targeting the Earth-Moon system and beyond that demand stringent positional accuracy and timing precision. Results, when available, will be reported in subsequent publications.
Acknowledgments
The work described here was carried out at the Jet Propulsion Laboratory, California Institute of Technology, Pasadena, California, under a contract with the National Aeronautics and Space Administration.
Appendix A Coordinate Transformation for the Moon System
A.1 Introducing the LCRS
To define the Lunicentric Coordinate Reference System (LCRS), we will adopt a similar approach to that used for the GCRS. Following the structure of the GCRS metric [19], the LCRS with coordinates ($\cal{T}$ = TCL, $\boldsymbol{\mathrm{\cal{X}}}$) is described by a metric of similar form. The components of this metric are tailored to account for the Moon’s gravitational field, tidal effects from nearby massive bodies, and the relativistic corrections required for high-precision trajectory modeling can be expressed in the form (see details in [9]):
$ds^{2}_{\tt LCRS}=\Big\{1-2c^{-2}\Big(U_{\tt M}({\cal T},\boldsymbol{\mathrm{\cal X}})+U^{\star}_{\rm tid}({\cal T},\boldsymbol{\mathrm{\cal X}})\Big)+{\cal O}(c^{-4})\Big\}c^{2}d{\cal T}^{2}-\Big\{1+2c^{-2}\Big(U_{\tt M}({\cal T},\boldsymbol{\mathrm{\cal X}})+U^{\star}_{\rm tid}({\cal T},\boldsymbol{\mathrm{\cal X}})\Big)+{\cal O}(c^{-4})\Big\}d\boldsymbol{\mathrm{\cal X}}^{2},$ (84)
where $U_{\tt M}({\cal T},\boldsymbol{\mathrm{\cal{X}}})$ represents the Newtonian gravitational potential of the isolated Moon, and $U^{\star}_{\rm tid}({\cal T},\boldsymbol{\mathrm{\cal{X}}})$ denotes the tidal potential generated by all other solar system bodies (excluding the Moon), both evaluated at the origin of the LCRS. The omitted terms, arising from $U_{\tt M}^{2}$, may contribute up to $\simeq 8.43\times 10^{-19}$, which is negligible for our purposes.
Analogous to the coordinate transformations between GCRS and BCRS described earlier, the time and position transformations between the LCRS ($\cal{T}$ = TCL, $\boldsymbol{\mathrm{\cal{X}}}$) and the BCRS ($t$ = TCB, $\boldsymbol{\mathrm{x}}$) can be expressed as follows:
${\cal T}$ $=$ $t-c^{-2}\Big\{\int^{t}_{t_{0}}\Big({\textstyle\frac{1}{2}}v_{\tt M}^{2}+\sum_{{\tt B}\not={\tt M}}{GM_{\tt B}\over r_{\tt BM}}\Big)dt+(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\boldsymbol{\mathrm{r}}_{\tt M})\Big\}+{\cal O}\Big(1.10\times 10^{-16}\Big)(t-t_{0}),$ (85)
$\boldsymbol{\mathrm{\cal X}}$ $=$ $\Big(1+c^{-2}\sum_{{\tt B}\not={\tt M}}{GM_{\tt B}\over r_{\tt BM}}\Big)\boldsymbol{\mathrm{r}}_{\tt M}+c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\boldsymbol{\mathrm{r}}_{\tt M})\boldsymbol{\mathrm{v}}_{\tt M}+(\boldsymbol{\mathrm{a}}_{\tt M}\cdot\boldsymbol{\mathrm{r}}_{\tt M})\boldsymbol{\mathrm{r}}_{\tt M}-{\textstyle\frac{1}{2}}r^{2}_{\tt M}\boldsymbol{\mathrm{a}}_{\tt M}\Big\}+{\cal O}\Big(1.28\times 10^{-12}~{\rm m}\Big),$ (86)
where $\boldsymbol{\mathrm{r}}_{\tt M}\equiv\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}_{\tt M}(t)$ with $\boldsymbol{\mathrm{x}}_{\tt M}$ and $\boldsymbol{\mathrm{v}}_{\tt M}=d\boldsymbol{\mathrm{x}}_{\tt M}/dt$ being the Moon’s position and velocity vectors in the BCRS. The uncertainty in (85) is determined by the magnitude of the $c^{-4}$-term omitted in (85) that behaves as $\sim c^{-4}\{-\frac{1}{8}v_{\tt M}^{2}-\frac{3}{2}v^{2}_{\tt M}GM_{\tt S}/r_{\tt M}+{\textstyle\frac{1}{2}}(GM_{\tt S}/r_{\tt M})^{2}\}\lesssim-1.22\times 10^{-16}=-10.50$ ps/d, which was omitted. The uncertainty in (86) is set by the omitted contribution from the solar quadrupole moment. The inverse transformations to (85)–(86) are obtained by moving the $c^{-2}$-terms on the other size of the equations with appropriate sign changes.
Using (84) and (85)–(86), we will derive the new relativistic time scales applicable to the Moon.
A.2 Transformation between TL and TDB
Analogous to (7), the time transformation from Lunar Time (TL) at the Moon’s surface to Luni-centric Coordinate Time (TCL) involves a change in the rate of time (see nomenclature discussion in [9])
$\frac{d{\tt TCL}}{d{\tt TL}}=\frac{1}{1-L_{\tt L}}=1+\frac{L_{\tt L}}{1-L_{\tt L}}\qquad{\rm or}\qquad\frac{d{\tt TL}}{d{\tt TCL}}=1-L_{\tt L}.$ (87)
For the Moon, constant $L_{\tt L}$ is derived from its gravity and rotational dynamics as below (see discussion in [19])
$L_{\tt L}=\frac{1}{c^{2}}\Big\{\frac{GM_{\tt M}}{R_{\tt MQ}}\big(1+{\textstyle\frac{1}{2}}J_{2\tt M}\big)+{\textstyle\frac{1}{2}}R_{\tt MQ}^{2}\omega_{\tt M}^{2}\Big\}+{\cal O}\big(3.27\times 10^{-15}\big).$ (88)
[…]
The constant $L_{\tt L}$ allows for the scaling of coordinates and mass parameters to maintain the invariance of the speed of light and the equations of motion within the LCRS during the transformation from TCL to TL. Thus, rather than using the coordinate time ${\cal T}={\tt TCL}$, spatial coordinates $\boldsymbol{\mathrm{\cal X}}$, and mass factors $(GM)_{\tt TCL}$ associated with the LCRS, we will apply the following scaling for the relevant quantities on the lunar surface:
${\tt TL}={\tt TCL}-L_{\tt L}({\tt TCL}-{\tt T}_{\tt L0}),\qquad\boldsymbol{\mathrm{\cal X}}_{\tt TL}=(1-L_{\tt L})\boldsymbol{\mathrm{\cal X}}_{\tt TCL},\qquad(GM)_{\tt TL}=(1-L_{\tt L})(GM)_{\tt TCL},$ (89)
where $\tt T_{L0}$} is the initial lunar time, which we will leave unspecified for the time being.
The transformation from TCL to TCB is determined using (85):
${\tt TCB}-{\tt TCL}=\Big\{\frac{1}{c^{2}}\int\Big({\textstyle\frac{1}{2}}v_{\tt M}^{2}+\sum_{{\tt B}\not={\tt M}}{GM_{\tt B}\over r_{\tt BM}}\Big)dt+\frac{1}{c^{2}}\big(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\boldsymbol{\mathrm{r}}_{\tt SM}\big)\Big\}_{\tt TCB}+{\cal O}(c^{-4}),$ (90)
[…]
Eq. (90) provides the mean rate between $\tt TCL$} and $\tt TCB$} as follows:
$\Big<\frac{d{\tt TCL}}{d{\tt TCB}}\Big>=1-L_{\tt H},$ (91)
[…]
To express $\tt TCB$} via $\tt TL$}, we introduce a constant $L_{\tt M}$, which defines the rate between $\tt TL$} and $\tt TCB$}. In analogy with the introduction of $L_{\tt B}$, constant $L_{\tt M}$ can be formally introduced as follows
$\Big<\frac{d{\tt TL}}{d{\tt TCB}}\Big>=1-L_{\tt M}.$ (92)
Then, by using the chain of time derivatives, we can establish the relationships between the constants $L_{\tt M},L_{\tt L}$ and $L_{\tt H}$, similar to (10). Following this approach, and with the use of (87), (91) and (92), we obtain the following:
$\Big<\frac{d{\tt TL}}{d{\tt TCB}}\Big>=\Big(\frac{d{\tt TL}}{d{\tt TCL}}\Big)\Big<\frac{d{\tt TCL}}{d{\tt TCB}}\Big>\qquad\Rightarrow\qquad(1-L_{\tt M})=(1-L_{\tt L})(1-L_{\tt H}),$ (93)
from this, the constant $L_{\tt M}$ is determined as $L_{\tt M}\simeq L_{\tt L}+L_{\tt H}-L_{\tt L}L_{\tt H}=1.485\,675\,294\times 10^{-8}\approx 1.283\,62$ ms/d.
To develop the $({\tt TCB}-{\tt TL})$ transformation, we apply (87), (93), and (90), leading to the following result:
${\tt TCB}-{\tt TL}=\frac{L_{\tt L}}{1-L_{\tt L}}({\tt TL}-{\tt T}_{\tt L0})+\frac{1}{1-L_{\tt M}}\Big\{\frac{1}{c^{2}}\int_{\tt T_{\tt L0}}^{\tt TL}\Big({\textstyle\frac{1}{2}}v_{\tt M}^{2}+\sum_{{\tt B}\not={\tt M}}{GM_{\tt B}\over r_{\tt BM}}\Big)dt+\frac{1}{c^{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\boldsymbol{\mathrm{r}}_{\tt 0M})\Big\}_{\tt TL},$ (94)
where $\boldsymbol{\mathrm{r}}_{\tt 0M}$ is the position vector of a specific site on the lunar surface in the LCRS.
Expression (94) allows us to formulate $({\tt TDB}-{\tt TL})$ as a function of TL. To achieve this, we use (11), (93), and (94), to express ${\tt TDB}-{\tt TL}$ as a function of TL as shown below
${\tt TDB}-{\tt TL}$ $=$ ${\tt TDB}_{0}-\frac{L_{\tt B}-L_{\tt L}}{1-L_{\tt L}}\,\big({\tt TL}-{\tt T}_{0}\big)+L_{\tt B}({\tt T}_{0}-{\tt T}_{\tt L0})+$ (95)
$+$ $\frac{1-L_{\tt B}}{1-L_{\tt M}}\Big\{\frac{1}{c^{2}}\int_{\tt T_{\tt L0}}^{\tt TL}\Big({\textstyle\frac{1}{2}}v_{\tt M}^{2}+\sum_{{\tt B}\not={\tt M}}{GM_{\tt B}\over r_{\tt BM}}\Big)d{\tt TL}+\frac{1}{c^{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\boldsymbol{\mathrm{r_{\tt 0M}}}_{\tt TL})\Big\}+{\cal O}(c^{-4}).$
It is useful to express the difference $({\tt TDB}-{\tt{\tt TL}})$ as a function of TDB. This can be achieved by applying (11), (89), (90), and (93), resulting in the following expression (see details in [9, 19]):
${\tt TDB}-{\tt TL}$ $=$ $\frac{1-L_{\tt L}}{1-L_{\tt B}}\,{\tt TDB}_{0}-\frac{L_{\tt B}-L_{\tt L}}{1-L_{\tt B}}\big({\tt TDB}-{\tt T}_{0}\big)+L_{\tt L}({\tt T}_{0}-{\tt T}_{\tt L0})+$ (96)
$+$ $\frac{1-L_{\tt L}}{1-L_{\tt B}}\Big\{\frac{1}{c^{2}}\int_{{\tt T_{0}+TDB_{0}}}^{\tt TDB}\Big({\textstyle\frac{1}{2}}v_{\tt M}^{2}+\sum_{{\tt B}\not={\tt M}}{GM_{\tt B}\over r_{\tt BM}}\Big)d{\tt TDB}+\frac{1}{c^{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\boldsymbol{\mathrm{r_{\tt 0M}}}_{\tt TDB})\Big\}+{\cal O}(c^{-4}).$
A.3 Relationships between TL and TT
Using expression (12) together with (96) and considering the Moon-related terms, it is instructive to express the BCRS position vector between a body B and the Moon as $\boldsymbol{\mathrm{r}}_{\tt BM}=\boldsymbol{\mathrm{r}}_{\tt BE}+\boldsymbol{\mathrm{r}}_{\tt EM}$, where $\boldsymbol{\mathrm{r}}_{\tt BE}=\boldsymbol{\mathrm{x}}_{\tt E}-\boldsymbol{\mathrm{x}}_{\tt B}$ is the position vector from the body B to the Earth, and $\boldsymbol{\mathrm{r}}_{\tt EM}=\boldsymbol{\mathrm{x}}_{\tt M}-\boldsymbol{\mathrm{x}}_{\tt E}$ is the Earth-Moon relative position vector, also $r_{\tt BM}\equiv|\boldsymbol{\mathrm{x}}_{\tt BM}|$, $r_{\tt EM}\equiv|\boldsymbol{\mathrm{x}}_{\tt EM}|$. Also, by representing $\boldsymbol{\mathrm{v}}_{\tt M}=\boldsymbol{\mathrm{v}}_{\tt E}+\boldsymbol{\mathrm{v}}_{\tt EM}$, where $\boldsymbol{\mathrm{v}}_{\tt E}$ is the BCRS velocity of the Earth and $\boldsymbol{\mathrm{v}}_{\tt EM}$ is the Earth-Moon relative velocity, we present expression $(\tt TL-TT)$ as a function of TT in the following form [9, 19]
${\tt TL}-{\tt TT}$ $=$ $\frac{L_{\tt G}-L_{\tt L}}{1-L_{\tt B}}\big({\tt TT}-{\tt T_{0}}\big)-L_{\tt L}({\tt T}_{0}-{\tt T}_{\tt LO})$- (97)
- $\frac{1}{c^{2}}\int_{\tt T_{0}}^{\tt TT}\Big\{{\textstyle\frac{1}{2}}v_{\tt EM}^{2}+\frac{GM_{\tt E}-2GM_{\tt M}}{r_{\tt EM}}+W^{\tt S}_{\tt EM}\Big\}d{\tt TT}-\frac{1}{c^{2}}(\boldsymbol{\mathrm{v}}_{\tt EM}\cdot\boldsymbol{\mathrm{\cal X}}_{\tt TT})+{\cal O}\Big(c^{-5};4.76\times 10^{-19}({\tt TT}-{\tt TT}_{0})\Big),$~~~~~
where the solar tidal potential at the Moon $W^{\tt S}_{\tt EM}$ is given by
$W^{\tt S}_{\tt EM}$ $=$ $\sum_{\ell=2}^{3}\frac{GM_{\tt S}}{r_{\tt SE}}\Big(\frac{r_{\tt EM}}{r_{\tt SE}}\Big)^{\ell}P_{\ell}(\boldsymbol{\mathrm{n}}_{\tt SE}\cdot\boldsymbol{\mathrm{n}}_{\tt EM})+{\cal O}\Big(4.30\times 10^{-19}\Big),$ (98)
where we kept the solar octupole tidal term $\ell=3$. The magnitude of this term was estimated to be $\simeq 1.68\times 10^{-16}$, which is small, but large enough to be part of the model. The error bound here is set by the solar $\ell=4$ tidal contribution, evaluated to be $c^{-2}GM_{\tt S}/r_{\tt SE}(r_{\tt EM}/r_{\tt SE})^{4}\simeq 4.30\times 10^{-19}$.
We can present result (97) in a different form by introducing the constant $L_{\tt EM}$ and periodic terms $P_{\tt EM}$ as below
$\frac{1}{c^{2}}\Big\{{\textstyle\frac{1}{2}}v_{\tt EM}^{2}+\frac{GM_{\tt E}-2GM_{\tt M}}{r_{\tt EM}}+W^{\tt S}_{\tt EM}\Big\}=L_{\tt EM}+\dot{P}_{\tt EM}(t)+{\cal O}\big(c^{-5};4.76\times 10^{-19}\big),$ (99)
where the constant rate $L_{\tt EM}\simeq L_{\tt H}-L_{\tt C}$ and the periodic terms $\dot{P}_{\tt EM}(t)\simeq\dot{P}_{\tt H}(t)-\dot{P}(t)$ are given as below:
$L_{\tt EM}$ $=$ $\frac{1}{c^{2}}\Big\{\Big<{\textstyle\frac{1}{2}}v_{\tt EM}^{2}+\frac{GM_{\tt E}-2GM_{\tt M}}{r_{\tt EM}}\Big>+\big<W^{\tt S}_{\tt EM}\big$$\Big\}$, (100)
$\dot{P}_{\tt EM}(t)$ $=$ $\frac{1}{c^{2}}\Big\{{\textstyle\frac{1}{2}}v_{\tt EM}^{2}+\frac{GM_{\tt E}-2GM_{\tt M}}{r_{\tt EM}}-\Big<{\textstyle\frac{1}{2}}v_{\tt EM}^{2}+\frac{GM_{\tt E}-2GM_{\tt M}}{r_{\tt EM}}\Big>+W^{\tt S}_{\tt EM}-\big<W^{\tt S}_{\tt EM}\big$$\Big\}$. (101)
Result (99), together with (100) and (101), provides valuable insight into the structure of the constant term $L_{\tt EM}$ and the periodic terms $P_{\tt EM}(t)$. These expressions can be used to explicitly establish the structure of the series $P_{\tt EM}(t)$.
Finally, using (99) in (97), we express $({\tt TL}-{\tt TT})$ as a function of TT:
${\tt TL}-{\tt TT}=\frac{L_{\tt G}-L_{\tt L}-L_{\tt EM}}{1-L_{\tt B}}\big({\tt TT}-{\tt T_{0}}\big)$ $-L_{\tt L}({\tt T}_{0}-{\tt T}_{\tt L0})-\Big(P_{\tt EM}({\tt TT})-P_{\tt EM}({\tt T_{0}})\Big)-\frac{1}{c^{2}}(\boldsymbol{\mathrm{v}}_{\tt EM}\cdot\boldsymbol{\mathrm{\cal X}}_{\tt TT})\,+$
$+{\cal O}\Big(4.76\times 10^{-19}({\tt TT}-{\tt T}_{0})\Big),$ (102)
where $\boldsymbol{\mathrm{\cal X}}_{\tt TT}$ is the TT-compatible lunicentric position of the lunar clock.
Considering the $O(c^{-2})$ term in $L_{\tt EM}$ (100), we use Moon–Earth relative speed of $v_{\tt EM}\approx 1022$ m/s, so the kinematic dilation contributes $c^{-2}\,\big<\tfrac{1}{2}\,v_{\tt EM}^{2}\big>\simeq 5.81\times 10^{-12},$ well above our $5\times 10^{-18}$ cutoff. Taking $r_{\tt EM}$ to be the instantaneous Earth–Moon separation, the Newtonian monopole term at the Earth–Moon distance was estimated to contribute up to $c^{-2}(GM_{\tt E}-2GM_{\tt M})/r_{\tt EM}\simeq 1.13\times 10^{-11}$. The solar quadrupole tide $W^{\tt S}_{\tt EM}$ yields $c^{-2}\big<W^{\tt S}_{\tt EM}\big>\simeq c^{-2}\tfrac{1}{4}\,{GM_{\tt S}\,r_{\tt EM}^{2}}/{r_{\tt SE}^{3}}\simeq 1.63\times 10^{-14}.$ As a result, collecting all the contributions, the secular‐drift coefficient for the Earth–Moon system is
$L_{\tt EM}=1.709\,390\,6\times 10^{-11}=1.4769\ \mu\mathrm{s}/\mathrm{d}.$ (103)
[…]
$L_{\tt B}-L_{\tt M}\simeq L_{\tt G}-L_{\tt L}-L_{\tt EM}=\big(60.2146-2.7121-1.4769\big)~\mu{\rm s/d}=56.0256~\mu{\rm s/d}.$ (104)
[…]
Now we consider the periodic term $P_{\tt EM}$, see (101). From the vis–viva relation for the Moon’s motion about the Earth–Moon barycenter given as $v^{2}_{\tt EM}(r)=(GM_{\tt E}+GM_{\tt M})\big({2}/{r_{\tt EM}}-{1}/{a_{\tt EM}}\big)$, with $r_{\tt EM}=a_{\tt EM}\bigl(1-e_{\tt M}\cos E\bigr)$, with $e_{\tt M}=0.0549$ being the Moon’s orbital eccentricity With these quantities, the orbital part of integrand in (99) reads
$\frac{1}{c^{2}}\Big\{\tfrac{1}{2}\,v_{\tt EM}^{2}+\frac{GM_{\tt E}-2GM_{\tt M}}{r_{\tt EM}}\Big\}=\frac{1}{c^{2}}\Big\{\frac{2GM_{\tt E}-GM_{\tt M}}{r_{\tt EM}}-\frac{GM_{\tt E}+GM_{\tt M}}{2a_{\tt EM}}\Big\}.$
Using this result in (101) and expanding $r_{\tt EM}$ to first order in $e_{\tt M}$, we can write $r_{\tt EM}=a_{\tt EM}\big(1-e_{\tt M}\cos[\omega_{\tt M}(t-t_{0})]+{\cal O}(e^{2}_{\tt M})\big)$, where $\omega_{\tt M}$ is the Moon’s mean orbital angular rate, we have [9, 19]
$\delta\dot{P}_{\tt EM}(t)$ $=\frac{1}{c^{2}}\Big\{{\textstyle\frac{1}{2}}v_{\tt EM}^{2}+\frac{GM_{\tt E}-2GM_{\tt M}}{r_{\tt EM}}-\Big<{\textstyle\frac{1}{2}}v_{\tt EM}^{2}+\frac{GM_{\tt E}-2GM_{\tt M}}{r_{\tt EM}}\Big$$\Big\}$=
$=\frac{1}{c^{2}}\big(2GM_{\tt E}-GM_{\tt M}\big)\Big(\frac{1}{r_{\tt EM}}-\frac{1}{a_{\tt EM}}\Big)\simeq\frac{1}{c^{2}}\big(2GM_{\tt E}-GM_{\tt M}\big)\frac{e_{\tt M}}{a_{\tt EM}}\cos[\omega_{\tt M}(t-t_{0})]=$
$=1.259\,047\times 10^{-12}\cos[\omega_{\tt M}(t-t_{0})]=0.109\,\cos[\omega_{\tt M}(t-t_{0})]~\mu{\rm s/d}.$ (105)
Integrating this result in time gives
$\delta P_{\tt EM}(t)$ $\simeq-\frac{1}{c^{2}}\big(2GM_{\tt E}-GM_{\tt M}\big)\frac{e_{\tt M}}{a_{\tt EM}\omega_{\tt M}}\sin[\omega_{\tt M}(t-t_{0})]=-0.473\,\sin[\omega_{\tt M}(t-t_{0})]~\mu{\rm s}.$ (106)
[…]
A.4 Position transformations
To develop the position transformation for the Moon, we define $\boldsymbol{\mathrm{\cal X}}_{\tt TCL}$ as the position vector of a site on the lunar surface in the LCRS and $\boldsymbol{\mathrm{r}}_{\tt M}$} as the corresponding vector observed from the BCRS (denoted ${\boldsymbol{\mathrm{r}}_{\tt M}}_{\tt TCB}$). The goal is to express the relationship between $\boldsymbol{\mathrm{\cal X}}_{\tt TCL}$ and ${\boldsymbol{\mathrm{r}}_{\tt M}}_{\tt TCB}$ in terms of the TL and TDB timescales. To achieve this, we use(89) to express the $\boldsymbol{\mathrm{\cal X}}_{\tt TCL}$ via $\tt TL$} and (11) to relate ${\boldsymbol{\mathrm{r}}_{\tt M}}_{\tt TCB}$ to ${\boldsymbol{\mathrm{r}}_{\tt M}}_{\tt TDB}$. This process leads to the following result:
$\boldsymbol{\mathrm{\cal X}}_{\tt TL}$ $=$ $\Big(1+L_{\tt B}-L_{\tt L}+c^{-2}\sum_{{\tt B}\not={\tt M}}{GM_{\tt B}\over r_{\tt BM}}\Big){\boldsymbol{\mathrm{r}}_{\tt M}}_{\tt TDB}+c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\boldsymbol{\mathrm{r}}_{\tt M})\boldsymbol{\mathrm{v}}_{\tt M}\Big\}_{\tt TDB}+{\cal O}\Big(1.00\times 10^{-7}~{\rm m}\Big),$~~~~ (107)
where $\boldsymbol{\mathrm{r}}_{\tt M}\equiv\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}_{\tt M}(t)$ and similar to (86), the error is determined by the omitted acceleration-dependent terms. The inverse transformation to that in (107) is straightforward; it is performed by shifting the $c^{-2}$-terms to the other side of the equation with the opposite signs. Note that the two leading corrections—the $c^{-2}\!\sum_{\tt B\neq M}GM_{\tt B}/r_{\tt BM}$ scale term and the $\tfrac{1}{2}c^{-2}(\mathbf{v}_{\tt M}\!\cdot\!\mathbf{r}_{\tt M})\,\mathbf{v}_{\tt M}$ Lorentz term—mirror their Earth analogues in (2).
The first group of terms in (107) represents a scale change. With $L_{\tt B}-L_{\tt L}=1.547\,380\,65\times 10^{-8}\approx 1.336\,94$ ms/d and $c^{-2}\big<\sum_{\tt B\not=M}GM_{\tt B}/r_{\tt BM}\big>=0.988438\times 10^{-8}$, the total is $2.53582\times 10^{-8}$, corresponding to a scale change of $4.41$ cm for the Moon’s radius. The annual variation is $\pm 1.65\times 10^{-10}$, or $\pm 0.29$ mm. Additionally, the Earth’s and Moon’s annual variations will affect the surface-to-surface distance along the center-to-center line, resulting in an annual variation of $1.34~{\rm mm}|\cos l^{\prime}|$.
The last term in (107) accounts for the Lorentz contraction. For a site on the Moon’s surface, this contraction amounts to $8.7$ mm. In the direction of motion, the combined effect of the scale change and Lorentz contraction totals $53$ mm, while perpendicular to the motion, the scale change is $44$ mm. These adjustments are essential for accurately determining the positions of retroreflectors used in lunar laser ranging (LLR), see [23].
A.5 Expressions to transform position, velocity and acceleration
Based on the analysis above and considering realistic measurement uncertainties, we use the magnitudes of the terms for a low-lunar-orbit (LLO) orbiter as the basis of our recommendation. Using the same approach as in Sec. II.3, we present the expressions recommended for position, velocity, and acceleration transformations for the LCRS [^4] Note that these expressions are analogous to those developed for the GCRS in (36), (42), and (51), respectively.:
$(\boldsymbol{\mathrm{x}}-\boldsymbol{\mathrm{x}}_{\tt M})_{\tt TDB}$ $=$ $\boldsymbol{\mathrm{\cal X}}_{\tt TL}-(L_{\tt B}-L_{\tt L})\boldsymbol{\mathrm{\cal X}}_{\tt TL}-c^{-2}\Big\{{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\boldsymbol{\mathrm{\cal X}})\boldsymbol{\mathrm{v}}_{\tt M}+\gamma U_{\tt M,\rm ext}\,\boldsymbol{\mathrm{\cal X}}\Big\}_{\tt TL}+{\cal O}\Big(3.34\times 10^{-7}~{\rm m}\Big),$ (108)
$(\dot{\boldsymbol{\mathrm{x}}}-\dot{\boldsymbol{\mathrm{x}}}_{\tt M})_{\tt TDB}$ $=$ $\dot{\boldsymbol{\mathrm{\cal X}}}_{\tt TL}-c^{-2}\Big\{\Big({\textstyle\frac{1}{2}}v_{\tt M}^{2}+(1+\gamma)U_{\tt M,\rm ext}+\big(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{X}}}\big)\Big)\dot{\boldsymbol{\mathrm{\cal X}}}+{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{X}}})\boldsymbol{\mathrm{v}}_{\tt M}\Big\}_{\tt TL}+{\cal O}\Big(7.82\times 10^{-9}~{\rm m/s}\Big),$~~~~~~ (109)
$(\ddot{\boldsymbol{\mathrm{x}}}-\ddot{\boldsymbol{\mathrm{x}}}_{\tt M})_{\tt TDB}$ $=$ $\ddot{\boldsymbol{\mathrm{\cal X}}}_{\tt TL}+(L_{\tt B}-L_{\tt L})\ddot{\boldsymbol{\mathrm{\cal X}}}_{\tt TL}-c^{-2}\Big\{\Big({v_{\tt M}^{2}}+(2+\gamma)U_{\tt M,\rm ext}+3\big(\boldsymbol{\mathrm{a}}_{\tt M}\cdot\boldsymbol{\mathrm{\cal X}}\big)+2\big(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{\cal X}}}\big)\Big)\ddot{\boldsymbol{\mathrm{\cal X}}}+$ (110)
$+$ $\Big((\boldsymbol{\mathrm{v}}_{\tt M}\cdot\boldsymbol{\mathrm{a}}_{\tt M})+(1+2\gamma)\dot{U}_{\tt M,\rm ext}+4\big(\boldsymbol{\mathrm{a}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{\cal X}}}\big)+\big(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}}\big)\Big)\dot{\boldsymbol{\mathrm{\cal X}}}+\Big((\boldsymbol{\mathrm{a}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{X}}})+{\textstyle\frac{1}{2}}(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}})\Big)\boldsymbol{\mathrm{v}}_{\tt M}+$
$+$ $\Big((\boldsymbol{\mathrm{a}}_{\tt M}\cdot\boldsymbol{\mathrm{\cal X}})+(\boldsymbol{\mathrm{v}}_{\tt M}\cdot\dot{\boldsymbol{\mathrm{\cal X}}})-(\dot{\boldsymbol{\mathrm{\cal X}}}\cdot\dot{\boldsymbol{\mathrm{\cal X}}})-({\boldsymbol{\mathrm{\cal X}}}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}})\Big)\boldsymbol{\mathrm{a}}_{\tt M}+(\boldsymbol{\mathrm{a}}_{\tt M}\cdot\ddot{\boldsymbol{\mathrm{\cal X}}})\boldsymbol{\mathrm{\cal X}}\Big\}_{\tt TL}+{\cal O}\Big(3.99\times 10^{-15}~{\rm m/s}^{2}\Big),\$
where ${\boldsymbol{\mathrm{v}}}_{\tt M}=d{\boldsymbol{\mathrm{x}}}_{\tt M}/dt$, ${\boldsymbol{\mathrm{a}}}_{\tt M}=d^{2}{\boldsymbol{\mathrm{x}}}_{\tt M}/dt^{2}$, and $U_{\tt M,\rm ext}$ and $\dot{U}_{\tt M,\rm ext}$ are given by (19) and (30), correspondingly by replacing E with M, see [19]. Also, the error terms are determined by the magnitude of similar terms as in (36), (42), and (51).
[…]
Appendix B Reference transform test vectors (for code validation)
Table 5 provides a fixed self-consistent test set for code validation. States are referenced to DE440 (TDB argument) and ITRF2020-to-GCRS realized by IERS (2010), [5]. Use double precision (or higher) and report agreement to the thresholds quoted below each equation.
[…]
Item Input Output Notes
Time map ${\tt TT}=\text{{\tt TT}(UTC 2025-01-01 00:00:00)}$ ${\tt TDB}=\text{TT}+$1.3\text{\times}{10}^{-3}$\,\mathrm{s}+\text{periodic}$ Eq. (14)
[…]
MEO state $(\boldsymbol{\mathrm{X}},\dot{\boldsymbol{\mathrm{X}}})$ at GPS altitude $(\boldsymbol{\mathrm{r}}_{\tt E},\dot{\boldsymbol{\mathrm{r}}}_{\tt E})_{\tt TDB}$ Eqs. (42),(51)